### v1

# Barron Space Representations for Elliptic PDEs with Homogeneous Boundary Conditions

Abstract
We study the approximation complexity of high-dimensional second-order elliptic PDEs with homogeneous boundary conditions on the unit hypercube, within the framework of Barron spaces. Under the assumption that the coefficients belong to suitably defined Barron spaces, we prove that the solution can be efficiently approximated by two-layer neural networks, circumventing the curse of dimensionality. Our results demonstrate the expressive power of shallow networks in capturing high-dimensional PDE solutions under appropriate structural assumptions.
1 Introduction
High-dimensional partial differential equations (PDEs) arise in a wide range of applications, including physics, finance, and control. Classical numerical schemes, such as finite difference and finite element methods, become computationally intractable as the spatial dimension $d$ increases, due to the so-called curse of dimensionality (CoD): to achieve an accuracy of $\varepsilon$, the computational cost scales as $\mathcal{O}(\varepsilon^{-d})$.
To overcome this barrier, a variety of deep learning-based approaches have been developed in recent years [EHJ17, HJE18, SS18, EY18, HZE19, RPK19, HSN20, PSMF20, ZHL21, NZGK21, EHJ22, KT24, KT25]. As an example, the Deep Ritz method (DRM) proposed by E and Yu [EY18] is a powerful approach for solving PDEs that admit a variational formulation. In this method, the PDE solution is represented by a deep neural network, and the network parameters are determined by minimizing the associated energy functional, with the integrals evaluated via Monte Carlo sampling. Since only first-order derivatives of the network output are required, this approach is particularly efficient for PDEs involving higher-order derivatives. Additionally, the Physics-Informed Neural Networks (PINNs) developed by Raissi et al. [RPK19] embeds the PDE and boundary conditions into the loss function by penalizing violations at selected collocation points. While this framework applies to a broader class of PDEs, it typically requires computing higher-order derivatives of the network output, which can reduce efficiency for PDEs with high-order terms.
Despite their differences in formulation, both DRM and PINNs share the common goal of approximating PDE solutions using neural networks. In practice, the performance of these methods depends critically on several theoretical aspects, including the representational capacity of the chosen neural network architecture, the design and convergence of the training algorithm, and the generalization ability of the resulting model to unseen inputs [SDK20, LLW21, LL22, DRM24]. These considerations naturally give rise to three fundamental theoretical questions in deep learning-based methods:
  • (Approximation) What is the minimal architectural complexity (e.g., depth, width) of a neural network required to approximate a target function with a prescribed accuracy?
  • (Optimization) Given such a network architecture, how can we design training algorithms with provable convergence guarantees to an optimal or near-optimal solution?
  • (Generalization) Under what conditions does the trained network exhibit strong generalization performance, ensuring low error on unseen data?
In this work, we focus on the first question. A prominent framework for addressing this problem is based on Barron functions. In his seminal work [Bar93], Barron proved that any real-valued function on $\mathbb{R}^{d}$ satisfying
$\int_{\mathbb{R}^{d}}|\xi||\widehat{f}(\xi)|{\,\mathrm{d}}\xi<\infty,$ (1.1)
a condition now referred to as having finite Barron norm, can be approximated in the $L^{2}$-norm by a two-layer neural network of the form
$\frac{1}{k}\sum_{i=1}^{k}a_{i}\,\sigma(w_{i}^{\top}x+b_{i}),$ (1.2)
with an approximation error independent of the input dimension $d$. This result provides a rigorous theoretical foundation for the ability of neural networks to circumvent the CoD in high-dimensional function approximation.
Following [Bar93], the approximation of Barron functions on bounded domains, such as hypercubes, has been investigated in [KB18, EY18, EMW22, EW22a], where similar dimension-independent approximation results were established. These works collectively suggest that Barron functions, whether defined on the whole space or on bounded domains like hypercubes, can be effectively approximated by two-layer neural networks without suffering from CoD; that is, the required number of neurons does not grow exponentially with the input dimension. This observation inspires a new paradigm for the study of high-dimensional PDEs:
If the coefficients of a PDE are such that its solution can be well approximated by a Barron function without incurring CoD, then a two-layer neural network can be used to efficiently approximate the solution in high dimensions.
This perspective was first realized by Chen et al. [CLL21], who studied second-order elliptic PDEs on $\mathbb{R}^{d}$. For bounded domains, Marwah et al. [MLLR23] established related results for a broad class of PDEs on the unit hypercube with Dirichlet boundary conditions, under certain regularity assumptions on the coefficients. The approach in [CLL21], motivated by [MLR21], constructs iterative schemes that converge exponentially fast to the solution while maintaining control of the Barron norm through careful analysis at each iteration. A similar methodology was also employed in [MLLR23].
For second-order elliptic PDEs on $\mathbb{R}^{d}$, the Barron function-based approximation framework is essentially complete [CLL21]. On bounded domains, [MLLR23] established results for Dirichlet problems on the unit hypercube, but these require additional assumptions on the Barron coefficients (cf. Remark 2.15). In contrast, the case of second-order elliptic PDEs with Neumann boundary conditions on bounded domains remains largely unaddressed.
1.1 Our Contributions
We study the approximation of solutions to second-order elliptic PDEs on the $d$-dimensional unit hypercube $\Omega\coloneqq(0,1)^{d}\subseteq\mathbb{R}^{d}$. Specifically, we consider the following two prototypical elliptic problems with homogeneous boundary conditions:
  • The Dirichlet Problem:
    $\begin{cases}-\nabla\cdot A(x)\nabla u+c(x)u=f(x),&\text{in }\Omega,\\ u=0,&\text{on }\partial\Omega,\end{cases}$ (1.3)
  • The Neumann Problem:
    $\begin{cases}-\nabla\cdot A(x)\nabla u+c(x)u=f(x),&\text{in }\Omega,\\ A(x)\nabla u\cdot\nu=0,&\text{on }\partial\Omega,\end{cases}$ (1.4)
where $A(x)$ is a real matrix-valued function, and $\nu$ denotes the outward unit normal vector on $\partial\Omega$. Our main results are stated informally below; precise formulations are given in Theorems 2.13 and 2.17.
Theorem 1.1 (Main Results, Informal Version).
Suppose the coefficients $A(x)$, $c(x)$, and $f(x)$ in equations (1.3) and (1.4) are suitable Barron functions. Then the weak solution $u^{*}$ to each equation can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network of the form
$\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i})\quad\text{or}\quad c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i}),$
using at most $\mathcal{O}(d^{\,C|\log\varepsilon|})$ neurons, where $C>0$ depends only on the coefficients of the PDE.
We briefly outline the strategy below, with full details and precise notation provided in the subsequent section.
  1. Step 1.
    For each of the equations (1.3) and (1.4), we construct the corresponding Sobolev gradient flow:
    $u_{t+1}=u_{t}-\alpha D\mathcal{E}(u_{t}).$
    We show that this scheme converges exponentially fast to the solution $u^{*}$ in the $H^{1}(\Omega)$-norm:
    $\|u_{t}-u^{*}\|_{H^{1}(\Omega)}\lesssim\beta^{t}\quad\text{for all }t\geq 0,$
    for some $0<\beta<1$ depending only on the PDE coefficients. (see Equations (3.19) and (3.20)). In fact, the iterative schemes previously proposed for variational PDEs in [CLL21, MLLR23] can also be interpreted as special cases of the Sobolev gradient flow.
  2. Step 2.
    We derive a recursive estimate for the Barron norm of $u_{t+1}$ in terms of that of $u_{t}$, showing that it grows at most at rate $\mathcal{O}(d^{2})$ (see Theorem 3.9). This ensures that all iterates $u_{t}$ remain Barron functions.
  3. Step 3.
    Applying the Barron function approximation theory developed in [EMW22] (see Theorem 3.14), any Barron function $g$ can be approximated by a two-layer neural network with cosine activation, and the $H^{1}(\Omega)$-error is bounded by
    $\left\|\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i})-g\right\|_{H^{1}(\Omega)}\leq\frac{\|g\|_{\mathcal{B}^{2}(\Omega)}}{\sqrt{k}},$
    For the $\operatorname{ReLU}$ activation case, we generalize the strategies in [LLW21] (see Theorem 3.15) and obtain
    $\left\|c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i})-g\right\|_{H^{1}(\Omega)}\lesssim\frac{\|g\|_{{\mathcal{B}}^{2}(\Omega)}}{\sqrt{k}}.$
    These results imply that the weak solution can be efficiently approximated by a two-layer neural network with either cosine or $\operatorname{ReLU}$ activation, without suffering from the CoD.
1.2 Related Works
In the Barron function-based framework, beyond approximating the solution by a Barron function, one may also pose the following question:
If the coefficients of the PDE are sufficiently regular such that the solution itself belongs to the Barron class, then it can be approximated by a two-layer neural network without incurring the CoD.
This viewpoint is often referred to as the regularity problem. It was first explored in [LLW21, LL22, EW22b], where the authors investigated the Poisson equation and the Schrödinger equation on the unit hypercube with Neumann boundary conditions. Subsequently, Chen et al. [CLLZ23] implemented this perspective for the stationary Schrödinger equation on $\mathbb{R}^{d}$. Regularity results for the Hamilton–Jacobi–Bellman equation in the whole space were established in [FL25], and for the electronic Schrödinger equation in [Yse25].
Apart from the Barron function-based approach, another influential deep learning methodology was proposed by E et al. [EHJ17, HJE18], who reformulated certain parabolic PDEs as backward stochastic differential equations (BSDEs), and further interpreted these BSDEs as stochastic control problems. These problems, in turn, can be viewed as instances of model-based reinforcement learning. This SDE-based framework was later extended by Sirignano and Spiliopoulos [SS18]. We refer to [TTY22, Rai23, KT24, LVR24] for additional applications of SDE-based methods to high-dimensional PDEs.
Alternative deep learning-based approaches have also been developed, typically relying on strong structural assumptions on the coefficients of the underlying PDEs. For instance, [GH21, GHJvW23] investigate classes of parabolic and Poisson equations that admit stochastic representations, such as those derived from the Feynman–Kac formula. In contrast, [MLR21] proposes neural network approximation conditions on the coefficients of second-order elliptic PDEs.
1.3 Organization
This paper is structured as follows. In Section 2, we formulate the main result. Section 2.1 is devoted to establishing the existence and uniqueness of weak solutions to problems (1.3) and (1.4). In Section 2.2, we introduce the Barron function spaces, which play a central role in our analysis. Building upon these foundations, we present main theorems in Section 2.3. Its proof is given in Section 3, where we provide a detailed exposition of the three-step strategy, with each step elaborated in the corresponding subsection, as briefly outlined in Section 1.1.
2 Formulation and Main Results
In this section, we formulate the problem under consideration and present our main results. Section 2.1 specifies the assumptions that ensure the existence and uniqueness of the weak solution. In Section 2.2, we introduce the Barron spaces and establish a basic consequence of their definition. Finally, Section 2.3 presents the main results of this work.
Throughout the remainder of this work, we let $\Omega\coloneqq(0,1)^{d}\subseteq\mathbb{R}^{d}$ denote the $d$-dimensional unit hypercube. We abbreviate the coefficient functions $A(x)$, $c(x)$, and $f(x)$ by $A$, $c$, and $f$, respectively. The associated second-order elliptic operator is defined as
$\mathcal{L}u\coloneqq-\nabla\cdot A\nabla u+cu.$
2.1 Existence and Uniqueness of Weak Solutions
To ensure the existence and uniqueness of weak solutions to equations (1.3) and (1.4), we impose the following assumption on their coefficients:
Assumption 2.1.
  1. (1)
    The real matrix-valued function $A(x)=(A_{ij}(x))_{1\leq i,j\leq d}$ defined on $\Omega$ is symmetric and uniformly elliptic. By uniform ellipticity, we mean that there exists a constant $a_{\min}>0$ such that
    $v^{\top}A(x)v\geq a_{\min}\|v\|^{2}\quad\text{for all }x\in\Omega\text{ and }v\in\mathbb{R}^{d},$
    where $\|v\|$ denotes the standard Euclidean norm in $\mathbb{R}^{d}$. Moreover, the operator norm of $A(x)$, defined as
    $\|A(x)\|_{\mathrm{op}}\coloneqq\sup_{x\in\Omega}\sup_{v\in\mathbb{R}^{d}\setminus\{0\}}\frac{\|A(x)v\|}{\|v\|},$
    is finite, and we denote $a_{\max}\coloneqq\|A(x)\|_{\mathrm{op}}$.
  2. (2)
    The scalar coefficient function $c(x)$ defined on $\Omega$ satisfies
    $0<c_{\min}\leq c(x)\leq c_{\max}<\infty\quad\text{for all }x\in\Omega,$
  3. (3)
    The source term $f(x)$ belongs to $L^{2}(\Omega)$.
Remark 2.2.
It is straightforward to verify that the finiteness of the operator norm of $A(x)$ implies the finiteness of its $L^{\infty}(\Omega)$-norm,
$\|A(x)\|_{L^{\infty}(\Omega)}\coloneqq\sum_{i,j=1}^{d}\|A_{ij}(x)\|_{L^{\infty}(\Omega)},$
and consequently, each entry $A_{ij}(x)\in L^{\infty}(\Omega)$. However, the converse is not true in general. A counterexample is given by $A(x)=\operatorname{diag}\{\delta_{P}(x),\ldots,\delta_{P}(x)\}$, where $\delta_{P}(x)$ is defined on $\Omega$ and equals $\infty$ if $x=P\coloneqq(1/2,\ldots,1/2)$, and 1 otherwise.
We refer to [Eva10, Lie13] for a comprehensive treatment of weak solutions in Sobolev spaces, and provide a brief summary here for completeness. A function $u_{D}^{*}\in H_{0}^{1}(\Omega)$ is said to be a weak solution of the Dirichlet problem (1.3) if it satisfies
$\int_{\Omega}\nabla u_{D}^{*}\cdot A\nabla v+cu_{D}^{*}v-fv{\,\mathrm{d}}x=0\quad\text{for all }v\in H_{0}^{1}(\Omega),$ (2.1)
A function $u_{N}^{*}\in H^{1}(\Omega)$ is said to be a weak solution of the Neumann problem (1.4) if it satisfies
$\int_{\Omega}\nabla u_{N}^{*}\cdot A\nabla v+cu_{N}^{*}v-fv{\,\mathrm{d}}x=0\quad\text{for all }v\in H^{1}(\Omega).$ (2.2)
Assumption 2.1, together with Remark 2.2 and the Lax–Milgram theorem [Eva10, Theorem 6.2.1], ensures the existence and uniqueness of weak solutions $u_{D}^{*}\in H_{0}^{1}(\Omega)$ and $u_{N}^{*}\in H^{1}(\Omega)$ to (1.3) and (1.4), respectively.
Remark 2.3.
Although, as noted in Remark 2.2, the class of matrix-valued functions with finite $L^{\infty}$-norm is strictly larger than that with bounded operator norm, and the subsequent analysis remains theoretically valid under the $L^{\infty}$-norm (cf. Lemma 3.6 and Theorem 3.4), we choose to adopt the operator norm. This choice is primarily motivated by the desire to ensure that the constants appearing in our main results exhibit milder dependence on the spatial dimension $d$. For example, when $A(x)=I_{d}$ is the $d\times d$ identity matrix, we have
$\|A(x)\|_{\operatorname{op}}=1,\quad\|A(x)\|_{L^{\infty}(\Omega)}=d.$
This demonstrates that the operator norm remains dimension-independent, while the $L^{\infty}$-norm grows linearly with $d$, making the former a more stable and dimension-robust choice.
2.2 Barron Spaces
To define the four types of Barron functions considered in this paper, we begin by introducing in Theorem 2.4 four families of functions in $L^{2}(\Omega)$, which serve as building blocks for representing elements in this space. Some of these families form bases of $L^{2}(\Omega)$. The justification is deferred to the Appendix A.1.
Theorem 2.4.
Let $\mathbb{N}\coloneqq\{0,1,2,\dots\}$ denote the set of natural numbers. For $k=(k_{1},\dots,k_{d})\in\mathbb{N}^{d}$, $\omega\in\mathbb{Z}^{d}$, and $x=(x_{1},\dots,x_{d})\in\Omega$, we define the following four families of functions in $L^{2}(\Omega)$:
  1. (1)
    $\left\{e^{i\pi\omega^{\top}x}:\omega\in\mathbb{Z}^{d}\right\}$,
  2. (2)
    $\left\{S_{k}(x)=\prod_{i=1}^{d}\sin(\pi k_{i}x_{i}):k\in\mathbb{N}^{d},\ k_{i}\neq 0\text{ for all }i\right\}$,
  3. (3)
    $\left\{C_{k}(x)=\prod_{i=1}^{d}\cos(\pi k_{i}x_{i}):k\in\mathbb{N}^{d}\right\}$,
  4. (4)
    $\left\{M_{k,i,j}(x)=C_{k}(x)\cdot\frac{\sin(\pi k_{i}x_{i})\sin(\pi k_{j}x_{j})}{\cos(\pi k_{i}x_{i})\cos(\pi k_{j}x_{j})}:k\in\mathbb{N}^{d},\ k_{i},k_{j}\neq 0\right\}$, where $1\leq i\neq j\leq d$ are fixed. That is, each function in this family is constructed such that the $i$-th and $j$-th components involve cosine terms, while all remaining components involve sine terms.
Every function in $L^{2}(\Omega)$ admits an $L^{2}$-expansion in family (1) with coefficients in $\mathbb{C}$, though the representation may be non-unique (see Example A.1). Families (2)-(4) form orthogonal bases of $L^{2}(\Omega)$ with coefficients in $\mathbb{R}$.
Remark 2.5.
By an $L^{2}$-expansion of a function in $L^{2}(\Omega)$ (e.g., with respect to family (2)), we mean that there exists a sequence $\{a_{k}\in\mathbb{R}:k\in\mathbb{N}^{d}\}$ such that
$\lim_{N\to\infty}\left\|g(x)-\sum_{\|k\|\leq N}a_{k}S_{k}(x)\right\|_{L^{2}(\Omega)}=0,$
denoted by
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)\quad\text{in }L^{2}(\Omega).$
A collection forms a basis of $L^{2}(\Omega)$ if every $g\in L^{2}(\Omega)$ admits a unique expansion in this family.
Definition 2.6.
Suppose $g(x)$ is a real-valued function defined on $\Omega$ and belongs to $L^{2}(\Omega)$.
  1. (1)
    We say that $g$ is an $e$-Barron function of weight $n\geq 0$ if its weighted $e$-Barron norm
    $\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\coloneqq\inf\left\{\sum_{\omega\in\mathbb{Z}^{d}}|a_{\omega}|\left(1+\pi^{n}\|\omega\|^{n}\right):g(x)=\sum_{\omega\in\mathbb{Z}^{d}}a_{\omega}e^{i\pi\omega^{\top}x}\;\text{in }L^{2}(\Omega)\right\}$
    is finite. The space $\mathcal{B}_{e}^{n}(\Omega)$ is referred to as the $e$-Barron space of weight $n$.
  2. (2)
    We say that $g$ is an $s$-Barron function of weight $n\geq 0$ if it admits the sine basis expansion
    $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)\quad\text{in }L^{2}(\Omega),$
    with finite weighted $s$-Barron norm
    $\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}\coloneqq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right).$
    The space $\mathcal{B}_{s}^{n}(\Omega)$ is referred to as the $s$-Barron space of weight $n$.
  3. (3)
    We say that $g$ is a $c$-Barron function of weight $n\geq 0$ if it admits the cosine basis expansion
    $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}C_{k}(x)\quad\text{in }L^{2}(\Omega),$
    with finite weighted $c$-Barron norm
    $\|g\|_{\mathcal{B}_{c}^{n}(\Omega)}\coloneqq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right).$
    The space $\mathcal{B}_{c}^{n}(\Omega)$ is referred to as the $c$-Barron space of weight $n$.
  4. (4)
    We say that $g$ is an $(i,j)$-mixed Barron function of weight $n\geq 0$ (for fixed indices $i\neq j$) if it admits a mixed basis expansion
    $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}M_{k,i,j}(x)\quad\text{in }L^{2}(\Omega),$
    with finite weighted $(i,j)$-mixed Barron norm
    $\|g\|_{\mathcal{B}_{i,j}^{n}(\Omega)}\coloneqq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right).$
    The space $\mathcal{B}_{i,j}^{n}(\Omega)$ is referred to as the $(i,j)$-mixed Barron space of weight $n$.
Remark 2.7.
The original definition of the Barron norm for functions on $\mathbb{R}^{d}$ was introduced in [Bar93], as given in equation (1.1). In contemporary literature, the version of the Barron norm defined via the Fourier transform—used in, e.g., [SX22, MLLR23, FL25]—is commonly referred to as the spectral Barron norm. In contrast, definitions that avoid the use of the Fourier transform, such as those based on probability measures (e.g., [CLL21, EMW22, SX23]), are typically referred to simply as the Barron norm. Our formulation is inspired by [LLW21, LL22, MLLR23], where cosine basis functions were used in [LLW21, LL22] and exponential basis functions in [MLLR23]. In this work, we introduce the sine-based Barron norm and the $(i,j)$-mixed Barron norm. The sine basis arises naturally when considering functions that vanish on the boundary, while the mixed norm is introduced for certain technical reasons (cf. Lemma 3.12). We also note that if the coefficient matrix $A(x)$ is diagonal, then the mixed Barron norm is not needed. This observation already covers many important families of PDEs, such as the stationary reaction–diffusion equations and static Schrödinger equations.
Remark 2.8.
Once a function $g$ is identified as a Barron function, its expansion in terms of the basis functions from Theorem 2.4 agrees with $g$ almost everywhere. (For the exponential basis, this holds if the weighted sum is suitably chosen to converge.) Consider the sine basis as an example. Suppose $g\in\mathcal{B}_{s}^{n}(\Omega)$, with its $L^{2}$-expansion given by $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$. Then, by the estimate
$\sum_{k\in\mathbb{N}^{d}}|a_{k}S_{k}(x)|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right)=\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}<\infty,$ (2.3)
the series $\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$ converges absolutely and uniformly on $\Omega$ by the Weierstrass $M$-test, and thus defines a continuous function on $\Omega$. Consequently, it coincides with $g(x)$ almost everywhere on $\Omega$. For simplicity, we will henceforth write
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$
without further qualification.
Remark 2.9.
Once a function $g$ is identified as a Barron function, its Barron norm can be computed using any reordering of its basis expansion. To illustrate, consider the sine basis. Suppose $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$ is an $s$-Barron function of weight $n$, and let $\tau$ be any permutation of $\mathbb{N}^{d}$. Assume further that $g(x)$ admits the expansion
$g(x)=\sum_{k\in\mathbb{N}^{d}}b_{k}S_{\tau(k)}(x)\quad\text{in }L^{2}(\Omega).$
Then, by the uniqueness of the $L^{2}(\Omega)$ expansion, we have $b_{k}=a_{\tau(k)}$ for all $k\in\mathbb{N}^{d}$, and hence
$\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}=\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right)=\sum_{k\in\mathbb{N}^{d}}|a_{\tau(k)}|\left(1+\pi^{n}\|\tau(k)\|^{n}\right)=\sum_{k\in\mathbb{N}^{d}}|b_{k}|\left(1+\pi^{n}\|\tau(k)\|^{n}\right),$
where the second equality is justified by the absolute convergence of the series. This reordering invariance will be used in the sequel, including in the proofs of Proposition 3.10 and Lemma 3.12.
Note that if $g$ is an $s$-Barron function, then any expansion with respect to the sine basis can be directly rewritten in terms of the exponential basis:
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}}\frac{a_{k}}{(2i)^{d}}\prod_{j=1}^{d}\left(e^{i\pi k_{j}x_{j}}-e^{-i\pi k_{j}x_{j}}\right)=\sum_{k\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\pm\frac{a_{k}}{(2i)^{d}}e^{i\pi v_{m}^{\top}x},$
where each vector $v_{m}=(\pm k_{1},\dots,\pm k_{d})\in\mathbb{Z}^{d}$ corresponds to one of the $2^{d}$ sign combinations. By the reordering invariance of the Barron norm (Remark 2.9) and the definition of the $e$-Barron norm, we obtain
$\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\left|\frac{a_{k}}{(2i)^{d}}\right|\left(1+\pi^{n}\|v_{m}\|^{n}\right)=\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right)=\|g\|_{\mathcal{B}_{s}^{n}(\Omega)},$
where we used that $\|v_{m}\|=\|k\|$.
A similar argument applies to other types of Barron functions, yielding the following result.
Proposition 2.10.
Let $\widetilde{\mathcal{B}}^{n}(\Omega)$ denote any of the spaces $\mathcal{B}_{s}^{n}(\Omega)$, $\mathcal{B}_{c}^{n}(\Omega)$, or $\mathcal{B}_{i,j}^{n}(\Omega)$. If $g\in\widetilde{\mathcal{B}}^{n}(\Omega)$, then
$\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\leq\|g\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}.$
In particular, the following inclusions hold for all $n\geq 0$:
$\mathcal{B}_{s}^{n}(\Omega),\;\mathcal{B}_{c}^{n}(\Omega),\;\mathcal{B}_{i,j}^{n}(\Omega)\subseteq\mathcal{B}_{e}^{n}(\Omega).$
2.3 Main Results
Our main results are established under the following assumptions on the coefficients appearing in equations (1.3) and (1.4).
Assumption 2.11 (Coefficients for Dirichlet Problem).
  1. (1)
    The entries of the coefficient matrix $A(x)=(A_{ij}(x))_{1\leq i,j\leq d}$ satisfy the following: $A_{ii}(x)\in\mathcal{B}_{c}^{1}(\Omega)$ for all $i$, and $A_{ij}(x)\in\mathcal{B}_{i,j}^{1}(\Omega)$ for all $i\neq j$. We define
    $\ell_{A,D}\coloneqq\max\left\{\max_{1\leq i\leq d}\|A_{ii}\|_{\mathcal{B}_{c}^{1}(\Omega)},\ \max_{1\leq i\neq j\leq d}\|A_{ij}\|_{\mathcal{B}_{i,j}^{1}(\Omega)}\right\}.$
  2. (2)
    The scalar coefficient $c(x)$ lies in $\mathcal{B}_{c}^{2}(\Omega)$, and we denote its Barron norm by $\ell_{c}\coloneqq\|c\|_{\mathcal{B}_{c}^{2}(\Omega)}$.
  3. (3)
    The source function $f(x)$ belongs to $\mathcal{B}_{s}^{0}(\Omega)$, with norm $\ell_{f,D}\coloneqq\|f\|_{\mathcal{B}_{s}^{0}(\Omega)}$.
Assumption 2.12 (Coefficients for Neumann Problem).
  1. (1)
    The entries of the coefficient matrix $A(x)=(A_{ij}(x))_{1\leq i,j\leq d}$ satisfy the following: $A_{ii}(x)\in\mathcal{B}_{s}^{1}(\Omega)$ for all $i$, and $A_{ij}(x)\in\mathcal{B}_{i,j}^{1}(\Omega)$ for all $i\neq j$. We define
    $\ell_{A,N}\coloneqq\max\left\{\max_{1\leq i\leq d}\|A_{ii}\|_{\mathcal{B}_{s}^{1}(\Omega)},\ \max_{1\leq i\neq j\leq d}\|A_{ij}\|_{\mathcal{B}_{i,j}^{1}(\Omega)}\right\}.$
  2. (2)
    The scalar coefficient $c(x)$ lies in $\mathcal{B}_{c}^{2}(\Omega)$, and we denote its Barron norm by $\ell_{c}\coloneqq\|c\|_{\mathcal{B}_{c}^{2}(\Omega)}$.
  3. (3)
    The source function $f(x)$ belongs to $\mathcal{B}_{c}^{0}(\Omega)$, with norm $\ell_{f,N}\coloneqq\|f\|_{\mathcal{B}_{c}^{0}(\Omega)}$.
For consistency in the subsequent analysis, we equip the Hilbert space $H_{0}^{1}(\Omega)$ with the inner product inherited from the standard inner product on $H^{1}(\Omega)$,
$(u,v)_{H^{1}(\Omega)}\coloneqq\int_{\Omega}\nabla u\cdot\nabla v+uv\,\mathrm{d}x.$
Although the norm induced by this inner product differs from the usual one on $H_{0}^{1}(\Omega)$, which is based solely on the gradient term, they are equivalent by Poincaré’s inequality (cf. [Eva10, Theorem 5.6.3]). This choice has no effect on our results.
We now present the first main result of the paper. The precise choice of parameters is given in Remark 2.14, and the full proof is deferred to the next section.
Theorem 2.13 (Main Result 1, Formal Version).
  1. (1)
    Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ denote the weak solution to the Dirichlet problem (1.3). Suppose Assumptions 2.1 and 2.11 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $N_{k}(x)\coloneqq\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos\left(w_{i}^{\top}x+b_{i}\right),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,D}^{2}\left(p_{D}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{D}(d)-1\right)^{2}}\right\rceil,$
    such that $\|N_{k}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
  2. (2)
    Let $u_{N}^{*}\in H^{1}(\Omega)$ denote the weak solution to the Neumann problem (1.4). Suppose Assumptions 2.1 and 2.12 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $N_{k}(x)\coloneqq\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos\left(w_{i}^{\top}x+b_{i}\right),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,N}^{2}\left(p_{N}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{N}(d)-1\right)^{2}}\right\rceil,$
    such that $\|N_{k}-u_{N}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
In other words, in both cases, the solution can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network with cosine activation using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where the constant $C>0$ depends only on the coefficients of the PDE.
Remark 2.14 (Parameter Explanation).
The constants and parameters appearing in Theorem 2.13 are defined as follows. Their derivation and role in the analysis will be discussed in detail in the next section:
$p_{D}(d)=\frac{2+\pi}{2\pi}\alpha_{\ast}\ell_{A,D}d^{2}+\alpha_{\ast}\ell_{c}+1,\qquad p_{N}(d)=\frac{2+\pi}{2\pi}\alpha_{\ast}\ell_{A,N}d^{2}+\alpha_{\ast}\ell_{c}+1,$
$T=\left\lceil\frac{\log\|f\|_{H^{-1}(\Omega)}+|\log\left(\varepsilon\lambda_{\min}/2\right)|}{|\log\beta_{\ast}|}\right\rceil,$
$\alpha_{\ast}=\frac{\lambda_{\min}}{2\lambda_{\max}^{2}},\qquad\beta_{\ast}=\left(1-\frac{\lambda_{\min}^{2}}{4\lambda_{\max}^{2}}\right)^{1/2},$
$\lambda_{\min}=\min\{a_{\min},c_{\min}\},\qquad\lambda_{\max}=\max\{a_{\max},c_{\max}\}.$
Remark 2.15.
Our result is stronger and more general than those in [CLL21, MLLR23] in the following aspects:
  1. (1)
    We do not require the solution to the Dirichlet problem (1.3) to be the restriction of a solution defined on $\mathbb{R}^{d}$, as is needed in [CLL21];
  2. (2)
    The work [MLLR23] claims a comparable network size. However, their derivation relies on the assumption that the coefficients in the Barron expansion vanish except for finitely many terms—an assumption not needed in our analysis;
  3. (3)
    In addition to considering only the Dirichlet problem as in [CLL21, MLLR23], we also provide a result for the Neumann problem.
Remark 2.16.
The quantities $p_{D}(d)$ and $p_{N}(d)$ can be chosen slightly smaller, although they still scale like $d^{2}$, as suggested by the argument in the proof of Theorem 3.9 in the next section. However, we retain the current definitions to ensure clarity and simplicity of exposition.
The effectiveness of periodic activation functions such as $\cos$ and $\sin$ has been empirically demonstrated in recent works [SMB+20, MTS21, BHRZ22, RRG25]. By contrast, for the more commonly used $\operatorname{ReLU}$ activation, analogous approximation results follow from the strategies in [LLW21], and the corresponding proof is also deferred to the next section.
Theorem 2.17 (Main Result 2, Formal Version).
  1. (1)
    Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ denote the weak solution to the Dirichlet problem (1.3). Suppose Assumptions 2.1 and 2.11 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $R_{k}(x)\coloneqq c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i}),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,D}^{2}q(d)\left(p_{D}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{D}(d)-1\right)^{2}}\right\rceil,$
    such that $\|R_{k}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
  2. (2)
    Let $u_{N}^{*}\in H^{1}(\Omega)$ denote the weak solution to the Neumann problem (1.4). Suppose Assumptions 2.1 and 2.12 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $R_{k}(x)\coloneqq c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i}),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,N}^{2}q(d)\left(p_{N}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{N}(d)-1\right)^{2}}\right\rceil,$
    such that $\|R_{k}-u_{N}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
Here, $q(d)=256d^{2}+128d+4$, and all parameters are as specified in Remark 2.14. Thus, in both cases, the solution can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network with $\operatorname{ReLU}$ activation using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where the constant $C>0$ depends only on the coefficients of the PDE.
Remark 2.18.
A similar approximation result for $\operatorname{ReLU}$ activation was also given in [LLW21]. However, our result is stronger and more general in the following aspects:
  1. (1)
    Our Barron space is substantially larger than that in [LLW21]. Their Barron norm is defined in a way similar to our $c$-Barron norm, but it is based on the $\ell^{1}$-norm $\|k\|_{1}\coloneqq|k_{1}|+\cdots+|k_{d}|$ rather than the Euclidean norm $\|k\|$ used here;
  2. (2)
    We consider general second-order elliptic PDEs with both Dirichlet and Neumann boundary conditions, rather than only the Poisson equation and the stationary Schrödinger equation with Neumann boundary conditions.
3 Proof of the Main Results
The objective of this section is to prove Theorems 2.13 and 2.17, following the three-step outline presented in Section 1.1. Our argument builds upon the methodology developed in [CLL21, LLW21, MLLR23, CLLZ24].
3.1 Step 1: Sobolev Gradient Flow
3.1.1 The Construction
It is well known that the solutions to the Dirichlet problem (1.3) and the Neumann problem (1.4) in the Sobolev space setting can be characterized as minimizers of suitable energy functionals. This classical characterization is summarized in the following proposition.
Proposition 3.1 (Variational Characterization of Weak Solutions).
  1. (1)
    Suppose Assumption 2.1 holds. Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ be the weak solution to the Dirichlet problem (1.3). Then $u_{D}^{*}$ is the unique minimizer of the energy functional
    $u_{D}^{*}=\mathop{\operatorname{arg\,min}}_{u\in H_{0}^{1}(\Omega)}\mathcal{E}_{D}(u),\quad\text{where}\quad\mathcal{E}_{D}(u)\coloneqq\int_{\Omega}\frac{1}{2}\nabla u\cdot A\nabla u+\frac{1}{2}cu^{2}-fu\,\mathrm{d}x.$ (3.1)
  2. (2)
    Suppose Assumption 2.1 holds. Let $u_{N}^{*}\in H^{1}(\Omega)$ be the weak solution to the Neumann problem (1.4). Then $u_{N}^{*}$ is the unique minimizer of the energy functional
    $u_{N}^{*}=\mathop{\operatorname{arg\,min}}_{u\in H^{1}(\Omega)}\mathcal{E}_{N}(u),\quad\text{where}\quad\mathcal{E}_{N}(u)\coloneqq\int_{\Omega}\frac{1}{2}\nabla u\cdot A\nabla u+\frac{1}{2}cu^{2}-fu\,\mathrm{d}x.$ (3.2)
Proof.
We only prove (1), as the proof of (2) follows similarly. By Stampacchia’s theorem (see [Bre11, Theorem 5.6]), it suffices to verify that
$\int_{\Omega}\nabla u_{D}^{*}\cdot A\nabla(v-u_{D}^{*})\,\mathrm{d}x\geq\int_{\Omega}f(v-u_{D}^{*})\,\mathrm{d}x\quad\text{for all }v\in H_{0}^{1}(\Omega).$
This inequality follows directly from the definition of the weak solution given in (2.1). ∎
The classical gradient descent method iteratively updates an initial guess by moving in the direction of the negative gradient of the objective function, aiming to converge to a minimizer. In the context of the PDEs we consider, Proposition 3.1 asserts that the weak solution is the unique minimizer of the associated energy functional. This naturally motivates the application of a gradient descent-type approach to solve the PDE. However, in contrast to the finite-dimensional setting, where the gradient is readily defined via classical derivatives, the notion of a gradient is not immediately available for energy functionals defined on infinite-dimensional spaces.
To overcome this difficulty, we recall an alternative way to characterize gradients. For a real-valued continuously differentiable function $g\colon\Omega\to\mathbb{R}$, its gradient $\nabla g(x)$ is the unique function $\nabla g(x)\colon\Omega\to\mathbb{R}^{d}$ satisfying
[…]
where $(\cdot,\cdot)_{\mathbb{R}^{d}}$ denotes the standard inner product in $\mathbb{R}^{d}$.
In our setting, taking the energy functional $\mathcal{E}_{D}$ for the Dirichlet problem as an example, we note that the inner product, subtraction, and scalar multiplication in (3.3) all have natural counterparts in the Hilbert space $H_{0}^{1}(\Omega)$. This motivates us to define the gradient of $\mathcal{E}_{D}$ as the map
$D_{H_{0}^{1}}\mathcal{E}_{D}\colon H_{0}^{1}(\Omega)\longrightarrow H_{0}^{1}(\Omega)$
such that
$(D_{H_{0}^{1}}\mathcal{E}_{D}(u),v)_{H^{1}(\Omega)}=\lim_{t\to 0}\frac{\mathcal{E}_{D}(u+tv)-\mathcal{E}_{D}(u)}{t}\quad\text{for all }u,v\in H_{0}^{1}(\Omega).$ (3.4)
Substituting the expression of $\mathcal{E}_{D}(u)$ from (3.1) into (3.4), we obtain
$\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u),v\right)_{H^{1}(\Omega)}=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }u,v\in H_{0}^{1}(\Omega).$
Here we use the inner product on $H_{0}^{1}(\Omega)$ inherited from the standard inner product on $H^{1}(\Omega)$. The uniqueness of $D_{H_{0}^{1}}\mathcal{E}_{D}$ follows from the Riesz representation theorem in the Hilbert space $H_{0}^{1}(\Omega)$, applied to the linear functional
$H_{0}^{1}(\Omega)\longrightarrow\mathbb{R},\quad v\longmapsto\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,$
for each $u\in H_{0}^{1}(\Omega)$.
A similar argument applies to the energy functional $\mathcal{E}_{N}$ for the Neumann problem. We summarize these results in the following definition.
Definition 3.2.
  1. (1)
    There exists a unique operator $D_{H_{0}^{1}}\mathcal{E}_{D}\colon H_{0}^{1}(\Omega)\to H_{0}^{1}(\Omega)$ such that
    $\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u),v\right)_{H^{1}(\Omega)}=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }u,v\in H_{0}^{1}(\Omega).$ (3.5)
    We call $D_{H_{0}^{1}}\mathcal{E}_{D}$ the Sobolev gradient of $\mathcal{E}_{D}$ with respect to $H_{0}^{1}(\Omega)$.
  2. (2)
    There exists a unique operator $D_{H^{1}}\mathcal{E}_{N}\colon H^{1}(\Omega)\to H^{1}(\Omega)$ such that
    $\left(D_{H^{1}}\mathcal{E}_{N}(u),v\right)_{H^{1}(\Omega)}=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }u,v\in H^{1}(\Omega).$ (3.6)
    We call $D_{H^{1}}\mathcal{E}_{N}$ the Sobolev gradient of $\mathcal{E}_{N}$ with respect to $H^{1}(\Omega)$.
Remark 3.3.
The concept of the Sobolev gradient has been widely employed in the study of PDEs, including applications to the Gross–Pitaevskii eigenvalue problem [HP20, Zha22, CLLZ24].
Analogous to the fact that the gradient of a function vanishes at its local minimum, the Sobolev gradient of an energy functional should also vanish at its minimizer. Indeed, by (2.1) and (2.2), we have
$\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u^{*}_{D}),v\right)_{H^{1}(\Omega)}$ $=0\quad\text{for all }v\in H_{0}^{1}(\Omega),$
$\left(D_{H^{1}}\mathcal{E}_{N}(u^{*}_{N}),v\right)_{H^{1}(\Omega)}$ $=0\quad\text{for all }v\in H^{1}(\Omega).$
By the non-degeneracy of the inner products on $H_{0}^{1}(\Omega)$ and $H^{1}(\Omega)$, it follows that
$D_{H_{0}^{1}}\mathcal{E}_{D}(u^{*}_{D})=D_{H^{1}}\mathcal{E}_{N}(u^{*}_{N})=0.$ (3.7)
In parallel with the classical gradient descent method, we define iterative schemes based on the Sobolev gradients $D_{H_{0}^{1}}\mathcal{E}_{D}$ and $D_{H^{1}}\mathcal{E}_{N}$ to approximate the solutions of the Dirichlet problem (1.3) and the Neumann problem (1.4), respectively. These schemes are referred to as Sobolev gradient flows, as they can be interpreted as discrete-time approximations of the corresponding continuous Sobolev gradient flows of the energy functionals.
  • Sobolev gradient flow for the Dirichlet problem:
    $u_{t+1}=u_{t}-\alpha D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}),$ (3.8)
    with initial condition $u_{0}=0$, where $\alpha>0$ is a step size to be specified later.
  • Sobolev gradient flow for the Neumann problem:
    $u_{t+1}=u_{t}-\alpha D_{H^{1}}\mathcal{E}_{N}(u_{t}),$ (3.9)
    with initial condition $u_{0}=0$, where $\alpha>0$ is a step size to be specified later.
3.1.2 Convergence Analysis
Theorem 3.4.
Assume Assumption 2.1 holds. The Sobolev gradient flow (3.8) for the Dirichlet problem (1.3) converges exponentially in the $H^{1}(\Omega)$-norm to the weak solution $u_{D}^{*}$. More precisely, there exists a constant $0<\beta<1$ such that
$\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\|u_{D}^{*}\|_{H^{1}(\Omega)}\beta^{t}\quad\text{for all }t\geq 0.$
An analogous result holds for the Neumann problem (1.4) under the Sobolev gradient flow (3.9), using the same choice of step size $\alpha$ and contraction factor $\beta$.
Remark 3.5.
Let
$\lambda_{\min}\coloneqq\min\{a_{\min},c_{\min}\},\quad\lambda_{\max}\coloneqq\max\{a_{\max},c_{\max}\},$
and define the step size and contraction factor by
$\alpha_{\ast}\coloneqq\frac{\lambda_{\min}}{2\lambda_{\max}^{2}},\qquad\beta_{\ast}\coloneqq\left(1-\frac{\lambda_{\min}^{2}}{4\lambda_{\max}^{2}}\right)^{1/2}.$ (3.10)
This choice of parameters is shown to be optimal for the Sobolev gradient flow (3.8) and (3.9) in the proof of Theorem 3.4 below. We therefore adopt this choice for all subsequent analysis after the proof.
The proof of Theorem 3.4 makes use of the following well-known lemma. For completeness, we include its proof in Appendix A.2.
Lemma 3.6.
Assume that the coefficient matrix $A(x)$ satisfies the condition in Assumption 2.1. Then, for all $u,v\in H^{1}(\Omega)$, we have
$\int_{\Omega}\nabla u\cdot A\nabla v\,{\,\mathrm{d}}x\leq a_{\max}\|\nabla u\|_{L^{2}(\Omega)}\|\nabla v\|_{L^{2}(\Omega)}.$
Proof of Theorem 3.4.
We provide the proof only for the Dirichlet problem (1.3), as the proof for the Neumann problem (1.4) follows analogously. Our goal is to establish that there exists a constant $0<\beta<1$ such that
$\|u_{t+1}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\beta\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}\quad\text{for all }t\geq 0.$
By the construction of the Sobolev gradient flow (3.8), we compute
$\|u_{t+1}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}$ $=\|u_{t}-u_{D}^{*}-\alpha D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}$
$=\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}+2\alpha\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}+\alpha^{2}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}.$ (3.11)
Estimating $(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}))_{H^{1}(\Omega)}$. Let $e_{t}\coloneqq u_{D}^{*}-u_{t}$. By the definition of the energy functional (3.1) and using (3.5), we have
$\mathcal{E}_{D}(u_{D}^{*})-\mathcal{E}_{D}(u_{t})$ $=\mathcal{E}_{D}(e_{t}+u_{t})-\mathcal{E}_{D}(u_{t})$
$=\int_{\Omega}\frac{1}{2}\nabla e_{t}\cdot A\nabla e_{t}+\frac{1}{2}ce_{t}^{2}{\,\mathrm{d}}x+\left(e_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}.$
Hence,
$\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}=\mathcal{E}_{D}(u_{D}^{*})-\mathcal{E}_{D}(u_{t})-\int_{\Omega}\frac{1}{2}\nabla e_{t}\cdot A\nabla e_{t}+\frac{1}{2}ce_{t}^{2}{\,\mathrm{d}}x.$ (3.12)
Since $u_{D}^{*}$ minimizes $\mathcal{E}_{D}$ over $H_{0}^{1}(\Omega)$ by Proposition 3.1, and $u_{t}\in H_{0}^{1}(\Omega)$ for all $t\geq 0$, it follows that $\mathcal{E}_{D}(u_{D}^{*})\leq\mathcal{E}_{D}(u_{t})$. Letting $\lambda_{\min}\coloneqq\min\{a_{\min},c_{\min}\}>0$, we deduce from (3.12) that
$\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}$ $\leq-\int_{\Omega}\frac{1}{2}\nabla e_{t}\cdot A\nabla e_{t}+\frac{1}{2}ce_{t}^{2}{\,\mathrm{d}}x$
$\leq-\frac{1}{2}\lambda_{\min}\int_{\Omega}\|\nabla e_{t}\|^{2}+\|e_{t}\|^{2}{\,\mathrm{d}}x=-\frac{1}{2}\lambda_{\min}\|e_{t}\|_{H^{1}(\Omega)}^{2}.$
Therefore,
$\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}\leq-\frac{1}{2}\lambda_{\min}\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}.$ (3.13)
Estimating $\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}$. By the vanishing property of the Sobolev gradient at the minimizer (3.7), we have
$\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{D}^{*}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}=0,\quad\text{for all }t\geq 0.$ (3.14)
Let $\lambda_{\max}\coloneqq\max\{a_{\max},c_{\max}\}$. Then, by combining (3.14), the definition of the Sobolev gradient in (3.5), Lemma 3.6, and the Cauchy–Schwarz inequality, we obtain
$\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}=\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}-\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{D}^{*}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}$
$=$ $\int_{\Omega}\nabla(u_{t}-u_{D}^{*})\cdot A\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})+c(u_{t}-u_{D}^{*})D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}){\,\mathrm{d}}x$
$\leq$ $a_{\max}\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}+c_{\max}\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}$
$\leq$ $\lambda_{\max}\left(\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}+\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}\right)$
$\leq$ $\lambda_{\max}\left(\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}^{2}+\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}^{2}\right)^{1/2}\left(\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}^{2}+\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}^{2}\right)^{1/2}$
$=$ $\lambda_{\max}\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}.$
This gives
$\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}\leq\lambda_{\max}\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}.$ (3.15)
Combining the estimates (3.11), (3.13), and (3.15), we obtain
$\|u_{t+1}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\left(1-\lambda_{\min}\alpha+\lambda_{\max}^{2}\alpha^{2}\right)^{1/2}\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}.$
Define the contraction factor
[…]
Then for sufficiently small $\alpha>0$, we can ensure that $0<\beta<1$, which completes the proof.
The optimality of the step size and contraction factor in (3.10) follows directly from analyzing the quadratic function in (3.16). ∎
Furthermore, by the definition of the weak solution to the Dirichlet problem in (2.1), the $H^{1}(\Omega)$-norm of $u_{D}^{*}$ can be bounded in terms of $\|f\|_{H^{-1}(\Omega)}$ as follows:
$\lambda_{\min}\|u_{D}^{*}\|_{H^{1}(\Omega)}^{2}\leq\int_{\Omega}\nabla u_{D}^{*}\cdot A\nabla u_{D}^{*}\,\mathrm{d}x+\int_{\Omega}c|u_{D}^{*}|^{2}\,\mathrm{d}x=\int_{\Omega}fu_{D}^{*}\,\mathrm{d}x\leq\|f\|_{H^{-1}(\Omega)}\|u_{D}^{*}\|_{H^{1}(\Omega)}.$ (3.17)
It follows that
$\|u_{D}^{*}\|_{H^{1}(\Omega)}\leq\frac{\|f\|_{H^{-1}(\Omega)}}{\lambda_{\min}}.$ (3.18)
Combining Theorem 3.4 with the bound in (3.18), we obtain
$\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\frac{\|f\|_{H^{-1}(\Omega)}}{\lambda_{\min}}\beta_{\ast}^{t}\quad\text{for all }t\geq 0.$ (3.19)
An analogous estimate holds for the weak solution $u_{N}^{*}$ to the Neumann problem under the Sobolev gradient flow (3.9):
$\|u_{t}-u_{N}^{*}\|_{H^{1}(\Omega)}\leq\frac{\|f\|_{H^{-1}(\Omega)}}{\lambda_{\min}}\beta_{\ast}^{t}\quad\text{for all }t\geq 0.$ (3.20)
Estimates (3.19) and (3.20) yield the following corollary.
Corollary 3.7.
Assume Assumption 2.1 holds. For any $\varepsilon\in(0,1/\lambda_{\min})$, let $\beta_{\ast}$ be defined as in (3.10). If
$T\geq\left\lceil\frac{\log\|f\|_{H^{-1}(\Omega)}+|\log(\varepsilon\lambda_{\min})|}{|\log\beta_{\ast}|}\right\rceil,$
then the iterates $u_{T}$ produced by both schemes (3.8) and (3.9) satisfy
$\|u_{T}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon,\quad\|u_{T}-u_{N}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
Remark 3.8.
In the estimate (3.17), we use the $H^{-1}(\Omega)$-norm of $f$ instead of the $L^{2}(\Omega)$-norm, since the $H^{-1}$-norm is smaller. This follows from the fact that the Poincaré constant $C_{P}$ for the unit hypercube $\Omega$ is $1/\pi\sqrt{d}$ (see Appendix A.3). That is,
$\|\phi\|_{L^{2}(\Omega)}\leq\frac{1}{\pi\sqrt{d}}\|\nabla\phi\|_{L^{2}(\Omega)}\quad\text{for all }\phi\in H_{0}^{1}(\Omega).$
Consequently, for every $f\in L^{2}(\Omega)$,
$\|f\|_{H^{-1}(\Omega)}$ $\coloneqq\sup\left\{\int_{\Omega}f\phi\,{\,\mathrm{d}}x:\phi\in H_{0}^{1}(\Omega),\;\|\nabla\phi\|_{L^{2}(\Omega)}\leq 1\right\}$
$\leq\sup\left\{\|f\|_{L^{2}(\Omega)}\|\phi\|_{L^{2}(\Omega)}:\phi\in H_{0}^{1}(\Omega),\;\|\nabla\phi\|_{L^{2}(\Omega)}\leq 1\right\}$
$\leq\sup\left\{\|f\|_{L^{2}(\Omega)}C_{P}\|\nabla\phi\|_{L^{2}(\Omega)}:\phi\in H_{0}^{1}(\Omega),\;\|\nabla\phi\|_{L^{2}(\Omega)}\leq 1\right\}$
$\leq\frac{1}{\pi\sqrt{d}}\|f\|_{L^{2}(\Omega)}.$
3.2 Step 2: Barron Norm Estimates
In this section, we focus on proving the following theorem.
Theorem 3.9.
Let $\alpha_{\ast}$ be defined as in (3.10), and set
[…]
  1. (1)
    Under Assumptions 2.1 and 2.11, the iterative scheme (3.8) for the Dirichlet problem (1.3) satisfies
    $\|u_{t+1}\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq p_{D}(d)\,\|u_{t}\|_{\mathcal{B}_{s}^{2}(\Omega)}+\frac{1}{2}\alpha_{\ast}\ell_{f,D}.$
  2. (2)
    Under Assumptions 2.1 and 2.12, the iterative scheme (3.9) for the Neumann problem (1.4) satisfies
    $\|u_{t+1}\|_{\mathcal{B}_{c}^{2}(\Omega)}\leq p_{N}(d)\,\|u_{t}\|_{\mathcal{B}_{c}^{2}(\Omega)}+\frac{1}{2}\alpha_{\ast}\ell_{f,N}.$
3.2.1 Sobolev Gradient Flow via the Inverse Elliptic Operator
For later use, we define the inverse elliptic operator $(I-\Delta)^{-1}$ under both Dirichlet and Neumann boundary conditions. Given any $\varphi\in L^{2}(\Omega)$, we define:
  • $(I-\Delta)_{D}^{-1}\varphi\in H_{0}^{1}(\Omega)$ denote the unique weak solution to
    $(I-\Delta)w=\varphi\quad\text{in }\Omega,\quad w=0\quad\text{on }\partial\Omega,$
  • $(I-\Delta)_{N}^{-1}\varphi\in H^{1}(\Omega)$ denote the unique weak solution to
    $(I-\Delta)w=\varphi\quad\text{in }\Omega,\quad\frac{\partial w}{\partial\nu}=0\quad\text{on }\partial\Omega.$
The existence and uniqueness of the weak solution in both cases are guaranteed by the Lax–Milgram theorem.
According to (3.5), we have
$\int_{\Omega}D_{H_{0}^{1}}\mathcal{E}_{D}(u)\,v+\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u)\cdot\nabla v\,\mathrm{d}x=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }v\in H_{0}^{1}(\Omega).$
Thus, $D_{H_{0}^{1}}\mathcal{E}_{D}(u)\in H_{0}^{1}(\Omega)$ is the unique weak solution to the elliptic problem
$(I-\Delta)w=\mathcal{L}u-f\quad\text{in }\Omega,\quad w=0\quad\text{on }\partial\Omega.$
This implies that the energy gradient for the Dirichlet problem can be expressed as
$D_{H_{0}^{1}}\mathcal{E}_{D}(u)=(I-\Delta)_{D}^{-1}(\mathcal{L}u-f).$
Hence, the iterative scheme (3.8) for the Dirichlet problem (1.3) becomes
$u_{t+1}=u_{t}-\alpha_{\ast}(I-\Delta)^{-1}_{D}(\mathcal{L}u_{t}-f).$ (3.22)
Similarly, from (3.6), we know that $D_{H^{1}}\mathcal{E}_{N}(u)\in H^{1}(\Omega)$ is the unique weak solution to
$(I-\Delta)w=\mathcal{L}u-f\quad\text{in }\Omega,\quad\frac{\partial w}{\partial\nu}=0\quad\text{on }\partial\Omega.$
Therefore, the energy gradient for the Neumann problem is given by
$D_{H^{1}}\mathcal{E}_{N}(u)=(I-\Delta)_{N}^{-1}(\mathcal{L}u-f).$
Hence, the iterative scheme (3.9) for the Neumann problem (1.4) becomes
$u_{t+1}=u_{t}-\alpha_{\ast}(I-\Delta)_{N}^{-1}(\mathcal{L}u_{t}-f).$ (3.23)
3.2.2 Analysis of the Barron Norm
A key ingredient in our analysis is the spectral property of the inverse elliptic operator. The result stated below will be used repeatedly in this section, and its proof follows from a direct computation.
Proposition 3.10.
For each $k\in\mathbb{N}^{d}$, we have
$(I-\Delta)_{D}^{-1}S_{k}(x)=\frac{1}{1+\pi^{2}\|k\|^{2}}S_{k}(x),\quad(I-\Delta)_{N}^{-1}C_{k}(x)=\frac{1}{1+\pi^{2}\|k\|^{2}}C_{k}(x).$
We next collect several algebraic and functional properties of Barron spaces that will be used in the proof of Theorem 3.9.
Proposition 3.11.
Let $\widetilde{\mathcal{B}}^{n}(\Omega)$ denote any of the spaces $\mathcal{B}_{e}^{n}(\Omega)$, $\mathcal{B}_{s}^{n}(\Omega)$, $\mathcal{B}_{c}^{n}(\Omega)$, or $\mathcal{B}_{i,j}^{n}(\Omega)$. Then:
  1. (1)
    (Linearity) For all $g_{1},g_{2}\in\widetilde{\mathcal{B}}^{n}(\Omega)$, and $\lambda_{1},\lambda_{2}\in\mathbb{R}$,
    $\|\lambda_{1}g_{1}+\lambda_{2}g_{2}\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}\leq|\lambda_{1}|\|g_{1}\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}+|\lambda_{2}|\|g_{2}\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}.$
  2. (2)
    (Monotonicity) Let $n\geq m\geq 0$. If $g\in\widetilde{\mathcal{B}}^{n}(\Omega)$, then
    $\|g\|_{\widetilde{\mathcal{B}}^{m}(\Omega)}\leq\|g\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}\quad\text{and}\quad\widetilde{\mathcal{B}}^{n}(\Omega)\subseteq\widetilde{\mathcal{B}}^{m}(\Omega).$
  3. (3)
    (Product estimates)
[…]
  4. (4)
    (Inverse operator estimates)
    $\left\|(I-\Delta)_{D}^{-1}g\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\frac{1}{2}\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}\quad\text{for }g\in\mathcal{B}_{s}^{0}(\Omega),$
    $\left\|(I-\Delta)_{N}^{-1}g\right\|_{\mathcal{B}_{c}^{2}(\Omega)}$ $=\frac{1}{2}\|g\|_{\mathcal{B}_{c}^{0}(\Omega)}\quad\text{for }g\in\mathcal{B}_{c}^{0}(\Omega).$
Proof.
Statements (1) and (2) follow directly from the definitions.
(3). We prove only the first inequality, as the second is analogous. Let
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x),\quad h(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}C_{k^{\prime}}(x).$
A similar estimation as in (2.3) shows that both series $\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)$ and $\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}C_{k^{\prime}}(x)$ converge absolutely and uniformly on $\Omega$. Hence, by Fubini’s theorem, the product $gh$ can be computed via termwise multiplication:
$gh$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\prod_{j=1}^{d}\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\cdot\frac{1}{2^{d}}\prod_{j=1}^{d}\left(\sin\left(\pi(k_{j}+k_{j}^{\prime})x_{j}\right)+\sin\left(\pi(k_{j}-k_{j}^{\prime})x_{j}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\pm\frac{1}{2^{d}}a_{k}b_{k^{\prime}}S_{v_{m}},$
where each $v_{m}=(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime})$ denotes one of the $4^{d}$ possible sign combinations such that $v_{m}\in\mathbb{N}^{d}$, which is ensured by the choice of signs in the last equality.
Using the fact that $\|v_{m}\|=\|\pm k\pm k^{\prime}\|\leq\|k\|+\|k^{\prime}\|$, we estimate
$1+\pi^{n}\|v_{m}\|^{n}\leq 1+\pi^{n}\left(\|k\|+\|k^{\prime}\|\right)^{n}\leq\left(1+\pi^{n}\|k\|^{n}\right)\left(1+\pi^{n}\|k^{\prime}\|^{n}\right).$ (3.24)
By the reordering invariance of the Barron norm stated in Remark 2.9 and estimate (3.24), we obtain
$\|gh\|_{\mathcal{B}_{s}^{n}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\pm\frac{1}{2^{d}}a_{k}b_{k^{\prime}}S_{v_{m}}(x)\right\|_{\mathcal{B}_{s}^{n}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{m=1}^{2^{d}}\frac{1}{2^{d}}|a_{k}b_{k^{\prime}}|\left(1+\pi^{n}\|v_{m}\|^{n}\right)$
$\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}|a_{k}||b_{k^{\prime}}|\left(1+\pi^{n}\|k\|^{n}\right)\left(1+\pi^{n}\|k^{\prime}\|^{n}\right)$
$=\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}\|h\|_{\mathcal{B}_{c}^{n}(\Omega)}.$
(4). We prove only the first identity, as the second follows analogously. Let
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x).$
Combining Proposition 3.10 with the following three estimates, valid for each $i=1,\dots,d$,
$\sum_{k\in\mathbb{N}^{d}}\left|a_{k}\frac{1}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\right|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\leq\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}<\infty,$
$\sum_{k\in\mathbb{N}^{d}}\left|a_{k}\frac{\pi k_{i}}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\frac{\cos(\pi k_{i}x)}{\sin(\pi k_{i}x)}\right|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\leq\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}<\infty,$
$\sum_{k\in\mathbb{N}^{d}}\left|a_{k}\frac{-\pi^{2}k_{i}^{2}}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\right|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\leq\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}<\infty,$
we conclude that the following series converge absolutely and uniformly on $\Omega$ by the Weierstrass $M$-test:
$\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x),\quad\partial_{i}\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x),\quad\partial_{ii}\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x).$
Therefore, we may freely exchange the order of summation and differentiation. In particular, we have
$(I-\Delta)_{D}\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x).$
Since the right-hand side belongs to $L^{2}(\Omega)$, the uniqueness of $(I-\Delta)_{D}^{-1}$ implies
$(I-\Delta)_{D}^{-1}\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x),$
and hence,
$(I-\Delta)_{D}^{-1}g=\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x).$
It now follows from Proposition 3.10 and the definition of the $s$-Barron norm that
$\left\|(I-\Delta)_{D}^{-1}g\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}=\left\|\sum_{k\in\mathbb{N}^{d}}a_{k}\frac{1}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}|a_{k}|=\frac{1}{2}\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}.$
Lemma 3.12.
Suppose the condition on $A(x)$ in Assumption 2.11 holds, and let $u\in\mathcal{B}_{s}^{2}(\Omega)$. Then for every $i,j=1,\dots,d$, we have
$\left\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ij}\partial_{j}u)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $\leq\frac{1}{\pi}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)},$ (3.25)
$\left\|(I-\Delta)^{-1}_{D}(A_{ij}\partial_{ij}u)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $\leq\frac{1}{2}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}.$ (3.26)
Under Assumption 2.12 and for $u\in\mathcal{B}_{c}^{2}(\Omega)$, the same estimates hold with $(I-\Delta)^{-1}_{N}$, $\ell_{A,N}$, and $\mathcal{B}_{c}^{2}(\Omega)$.
Proof.
We prove the case under Assumption 2.11; the case under Assumption 2.12 follows by the same argument.
Proof of (3.25) in the case $i=j$. Let
$A_{ii}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}C_{k}(x),\quad u(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}S_{k^{\prime}}(x).$
Since
$\sum_{k\in\mathbb{N}^{d}}|a_{k}\partial_{i}C_{k}(x)|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}\pi k_{i}|\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi\|k\|\right)=\|g\|_{\mathcal{B}_{c}^{1}(\Omega)}<\infty,$
it follows from the Weierstrass $M$-test that the series $\sum_{k\in\mathbb{N}^{d}}a_{k}\partial_{i}C_{k}(x)$ converges absolutely and uniformly on $\Omega$. Therefore, we are justified in interchanging the order of summation and differentiation. A similar argument applies to the series $\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}\,\partial_{i}S_{k^{\prime}}(x)$, which also converges absolutely and uniformly. Moreover, since both series are absolutely convergent, we may apply Fubini’s theorem to freely interchange the order of summation when taking products of their terms. Thus, we compute
$\partial_{i}A_{ii}\partial_{i}u$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\left(\partial_{i}C_{k}\partial_{i}S_{k^{\prime}}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)\cdot\sin\left(\pi k_{i}x_{i}\right)\cos\left(\pi k_{i}^{\prime}x_{i}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\cdot\frac{1}{2^{d}}\prod_{m=1}^{d}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)\pm\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}S_{v_{n}}(x),$
where each $v_{n}=\left(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime}\right)$, with appropriately chosen signs such that $v_{n}\in\mathbb{N}^{d}$.
Combining the reordering invariance of the Barron norm (Remark 2.9), a similar summation–differentiation interchange argument in the proof of Proposition 3.11 (4), and Proposition 3.10, we obtain
$\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ii}\partial_{i}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}(I-\Delta)^{-1}_{D}S_{v_{n}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\frac{1}{1+\pi^{2}\|v_{n}\|^{2}}S_{v_{n}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}k_{i}^{\prime}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k\|\|k^{\prime}\|^{2}$
$\leq\frac{1}{\pi}\|A_{ii}\|_{\mathcal{B}_{c}^{1}(\Omega)}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\frac{1}{\pi}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}.$
Proof of (3.25) in the case $i\neq j$. Suppose
$A_{ij}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}M_{k,i,j}(x),\quad u(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}S_{k^{\prime}}(x).$
By a similar argument of absolute and uniform convergence as in the proof of (3.25), the following computation is justified:
$\partial_{i}A_{ij}\partial_{j}u$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\quad\quad\cdot\cos\left(\pi k_{i}x_{i}\right)\sin\left(\pi k_{i}^{\prime}x_{i}\right)\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\cdot\frac{1}{2^{d}}\prod_{m=1}^{d}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)\pm\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}S_{v_{n}}(x),$
where each $v_{n}=\left(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime}\right)$, with appropriately chosen signs such that $v_{n}\in\mathbb{N}^{d}$.
Combining the reordering invariance of the Barron norm (Remark 2.9), a similar summation–differentiation interchange argument in the proof of Proposition 3.11 (4), and Proposition 3.10, we obtain
$\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ij}\partial_{j}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}(I-\Delta)^{-1}_{D}S_{v_{n}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\frac{1}{1+\pi^{2}\|v_{n}\|^{2}}S_{v_{n}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}k_{j}^{\prime}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k\|\|k^{\prime}\|^{2}$
$\leq\frac{1}{\pi}\|A_{ij}\|_{\mathcal{B}_{i,j}^{1}(\Omega)}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\frac{1}{\pi}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}.$
Proof of (3.26) in the case $i=j$. Let
$A_{ii}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}C_{k}(x),\quad u(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}S_{k^{\prime}}(x).$
By a similar argument of absolute and uniform convergence as in the proof of (3.25), the following computation is justified:
$A_{ii}\partial_{ii}u$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\left(C_{k}\partial_{ii}S_{k^{\prime}}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\prod_{m=1}^{d}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\cdot\frac{1}{2^{d}}\prod_{m=1}^{d}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)-\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}S_{v_{n}}(x),$
where each $v_{n}=\left(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime}\right)$, with appropriately chosen signs such that $v_{n}\in\mathbb{N}^{d}$.
Combining the reordering invariance of the Barron norm (Remark 2.9), a similar summation–differentiation interchange argument in the proof of Proposition 3.11 (4), and Proposition 3.10, we obtain
$\|(I-\Delta)^{-1}_{D}(A_{ii}\partial_{ii}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}(I-\Delta)^{-1}_{D}S_{v_{n}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\frac{1}{1+\pi^{2}\|v_{n}\|^{2}}S_{v_{n}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}^{\prime 2}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k^{\prime}\|^{2}$
$\leq\frac{1}{2}\|A_{ii}\|_{\mathcal{B}_{c}^{0}(\Omega)}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\frac{1}{2}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}.$
Proof of (3.26) in the case $i\neq j$. Suppose
$A_{ij}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}M_{k,i,j}(x),\quad u(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}S_{k^{\prime}}(x).$
By a similar argument of absolute and uniform convergence as in the proof of (3.25), the following computation is justified:
$A_{ij}\partial_{ij}u$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\quad\quad\cdot\sin\left(\pi k_{i}x_{i}\right)\cos\left(\pi k_{i}^{\prime}x_{i}\right)\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\cdot\frac{1}{2^{d}}\prod_{m=1}^{d}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)\pm\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}S_{v_{n}}(x),$
where each $v_{n}=\left(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime}\right)$, with appropriately chosen signs such that $v_{n}\in\mathbb{N}^{d}$.
Combining the reordering invariance of the Barron norm (Remark 2.9), a similar summation–differentiation interchange argument in the proof of Proposition 3.11 (4), and Proposition 3.10, we obtain
$\|(I-\Delta)^{-1}_{D}(A_{ij}\partial_{ij}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}(I-\Delta)^{-1}_{D}S_{v_{n}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\frac{1}{1+\pi^{2}\|v_{n}\|^{2}}S_{v_{n}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}^{\prime}k_{j}^{\prime}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k^{\prime}\|^{2}$
$\leq\frac{1}{2}\|A_{ij}\|_{\mathcal{B}_{i,j}^{0}(\Omega)}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\frac{1}{2}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}$
Lemma 3.13.
  1. (1)
    Suppose the condition on $c(x)$ in Assumption 2.11 holds, and let $u\in\mathcal{B}_{s}^{2}(\Omega)$. Then we have
    $\|(I-\Delta)^{-1}_{D}(cu)\|_{{\mathcal{B}}_{s}^{2}(\Omega)}\leq\ell_{c}\|u\|_{{\mathcal{B}}_{s}^{2}(\Omega)}.$
  2. (2)
    Suppose the condition on $c(x)$ in Assumption 2.12 holds, and let $u\in{\mathcal{B}}_{c}^{2}(\Omega)$. Then we have
    $\|(I-\Delta)^{-1}_{N}(cu)\|_{{\mathcal{B}}_{c}^{2}(\Omega)}\leq\ell_{c}\|u\|_{{\mathcal{B}}_{c}^{2}(\Omega)}.$
Proof.
We provide the proof of (1), as the proof of (2) follows by the same argument. By part (3) of Proposition 3.11, we have $cu\in\mathcal{B}_{s}^{2}(\Omega)$. It then follows from part (4) of the same proposition that
$\|(I-\Delta)^{-1}_{D}(cu)\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\|cu\|_{\mathcal{B}_{s}^{2}(\Omega)}.$
Applying part (3) again yields the desired result. ∎
Proof of Theorem 3.9.
Note that
$\mathcal{L}u-f=-\nabla\cdot A\nabla u+cu-f=-\sum_{i,j=1}^{d}\partial_{i}A_{ij}\partial_{j}u-\sum_{i,j=1}^{d}A_{ij}\partial_{ij}u+cu-f.$
The result follows from the linearity of $(I-\Delta)^{-1}_{D}$ and $(I-\Delta)^{-1}_{N}$, together with (3.22), (3.23), Proposition 3.11, and Lemmas 3.12 and 3.13. ∎
3.3 Step 3: Neural Network Approximation
The following approximation theorem for Barron functions by neural networks is analogous to [EMW22, Theorem 4], and our proof strategy is a refinement of their argument. To the best of our knowledge, this remains the standard approach in the literature; similar strategies are employed in the approximation results for Barron functions in [CLL21, CLLZ23, FL25].
Theorem 3.14.
Let $g\in\mathcal{B}_{e}^{2}(\Omega)$ be an $e$-Barron function. Suppose that the activation function $\sigma$ in the two-layer neural network of the form (1.2) is the cosine function. Then, for any fixed positive integer $k$, there exists a set of parameters
$\left\{(a_{i},w_{i},b_{i})\in\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right\}_{1\leq i\leq k}$
such that
$\left\|\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma\left(w_{i}^{\top}x+b_{i}\right)-g(x)\right\|_{H^{1}(\Omega)}\leq\frac{\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}}{\sqrt{k}}.$ (3.27)
Proof.
Let $\varepsilon>0$ be fixed. Suppose
$g(x)=\sum_{\omega\in\mathbb{Z}^{d}}c_{\omega}e^{i\pi\omega^{\top}x}=\sum_{\omega\in\mathbb{Z}^{d}}c_{\omega}\cos\left(\pi\omega^{\top}x\right)+ic_{\omega}\sin\left(\pi\omega^{\top}x\right)$
satisfies
$\sum_{\omega\in\mathbb{Z}^{d}}|c_{\omega}|\left(1+\pi^{2}\|\omega\|^{2}\right)\leq\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}+\varepsilon.$ (3.28)
We write $c_{\omega}=c_{\omega,R}+ic_{\omega,I}$ with $c_{\omega,R},c_{\omega,I}\in\mathbb{R}$. Since $g(x)$ is real-valued, we have
$g(x)=\operatorname{Re}(g(x))=\sum_{\omega\in\mathbb{Z}^{d}}-c_{\omega,I}\sin\left(\pi\omega^{\top}x\right)+c_{\omega,R}\cos\left(\pi\omega^{\top}x\right).$ (3.29)
Let $\operatorname{sgn}(x)$ denote the sign function, which equals $1$ if $x>0$, 0 if $x=0$, and 1-1 if $x<0$. Define
[…]
Then, using the standard cosine addition identity, we obtain
$c_{\omega,I}\sin\left(\pi\omega^{\top}x\right)+c_{\omega,R}\cos\left(\pi\omega^{\top}x\right)=|c_{\omega}|\cos\left(\pi\omega^{\top}x+\theta_{\omega}\right),$ (3.30)
where $|c_{\omega}|=\sqrt{c_{\omega,R}^{2}+c_{\omega,I}^{2}}$. Combining (3.30) with (3.29), we obtain
$g(x)=\sum_{\omega\in\mathbb{Z}^{d}}|c_{\omega}|\cos\left(\pi\omega^{\top}x+\theta_{\omega}\right).$ (3.31)
By (3.28), we know that
[…]
Thus, the measure
[…]
defines a probability measure on $\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}$, where $\delta_{(Z,\pi\omega,\theta_{\omega})}$ denotes the Dirac measure on $\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}$. We equip the space $(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}$, endowed with its Borel $\sigma$-algebra, with the product probability measure ${\mu(a,w,b)}^{\otimes k}$. Let
[…]
and define the pointwise approximation error between the neural network and the target function $g$ by
[…]
If we regard $\Theta$ as a random variable on the probability space $\left((\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k},{\mu(a,w,b)}^{\otimes k}\right)$, then to prove the theorem, it suffices to establish that
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{H^{1}(\Omega)}\leq\frac{\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}^{2}}{k}.$ (3.32)
Observe that the expected $H^{1}$-error decomposes as
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{H^{1}(\Omega)}=\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}+\mathbb{E}_{\mu^{\otimes k}}\|\nabla R(\Theta;x)\|_{L^{2}(\Omega)}^{2},$ (3.33)
We proceed to estimate the two terms on the right-hand side separately.
Expectation of $\|R(\Theta;x)\|_{L^{2}(\Omega)}^{2}$. Since we assume that $\sigma$ is the cosine function, it follows from (3.31) that
$\mathbb{E}_{\mu(a,w,b)}\left[a\sigma(w^{\top}x+b)\right]$ $=\int_{\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}}a\cos\left(w^{\top}x+b\right){\,\mathrm{d}}\mu(a,w,b)$
$=\sum_{\omega\in\mathbb{Z}^{d}}\frac{|c_{\omega}|}{Z}Z\cos\left(\pi\omega^{\top}x+\theta_{\omega}\right)=g(x).$ (3.34)
It follows from (3.34) that for every $1\leq i\neq j\leq k$, we have
$\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)\left(a_{j}\sigma(w_{j}^{\top}x+b_{j})-g(x)\right){\,\mathrm{d}}{\mu(a,w,b)}^{\otimes k}=0.$ (3.35)
Thus, combining (3.35) and (3.34), we compute that
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$ $=\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\int_{\Omega}\left(\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)^{2}{\,\mathrm{d}}x\mathrm{d}{\mu(a,w,b)}^{\otimes k}$
$=\frac{1}{k^{2}}\int_{\Omega}\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(\sum_{i=1}^{k}\left(a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)\right)^{2}{\,\mathrm{d}}{\mu(a,w,b)}^{\otimes k}\mathrm{d}x$
$=\frac{1}{k^{2}}\int_{\Omega}\sum_{i=1}^{k}\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)^{2}{\,\mathrm{d}}{\mu(a,w,b)}^{\otimes k}\mathrm{d}x$
$=\frac{1}{k}\int_{\Omega}\int_{\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}}\left(a\sigma(w^{\top}x+b)-g(x)\right)^{2}{\,\mathrm{d}}\mu(a,w,b)\mathrm{d}x$
$=\frac{1}{k}\int_{\Omega}\int_{\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}}\left(a\sigma(w^{\top}x+b)-\mathbb{E}_{\mu(a,w,b)}\left[a\sigma(w^{\top}x+b)\right]\right)^{2}{\,\mathrm{d}}\mu(a,w,b)\mathrm{d}x$
$=\frac{1}{k}\int_{\Omega}\operatorname{Var}_{\mu(a,w,b)}\left(a\sigma(w^{\top}x+b)\right){\,\mathrm{d}}x.$
Since $\sigma$ is the cosine function, we have
$\operatorname{Var}_{\mu(a,w,b)}\left(a\sigma(w^{\top}x+b)\right)\leq\mathbb{E}_{\mu(a,w,b)}\left[\left(a\sigma(w^{\top}x+b)\right)^{2}\right]\leq\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\right].$
Thus,
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}\leq\frac{m(\Omega)}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\right]=\frac{1}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\right].$ (3.36)
where $m(\Omega)$ denotes the Lebesgue measure of $\Omega$, which equals $1$ since $\Omega=(0,1)^{d}$.
Expectation of $\|\nabla R(\Theta;x)\|_{L^{2}(\Omega)}^{2}$. Let $\langle w_{i},e_{j}\rangle$ denote the standard Euclidean inner product of $w_{i}$ and $e_{j}$ in $\mathbb{R}^{d}$; that is, $\langle w_{i},e_{j}\rangle$ is the $j$-th component of $w_{i}\in\mathbb{R}^{d}$. Differentiating both sides of (3.34) with respect to $x_{j}$, it is straightforward to verify that for every $1\leq i\neq\ell\leq k$ and for each $1\leq j\leq d$, we have
$\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(a_{i}\langle w_{i},e_{j}\rangle\sigma^{\prime}(w_{i}^{\top}x+b_{i})-\partial_{j}g(x)\right)\left(a_{\ell}\langle w_{\ell},e_{j}\rangle\sigma^{\prime}(w_{\ell}^{\top}x+b_{\ell})-\partial_{j}g(x)\right){\,\mathrm{d}}\mu^{\otimes k}=0,$
$\mathbb{E}_{(a,w,b)}\left[a\left\langle w,e_{j}\right\rangle\sigma^{\prime}(w^{\top}x+b)\right]=\partial_{j}g(x).$
Thus, for every $1\leq j\leq d$, we have
$\mathbb{E}_{\mu^{\otimes k}}\left\|\partial_{j}R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$ $=\int_{\left(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right)^{k}}\int_{\Omega}\left(\partial_{j}\left(\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma\left(w_{i}^{\top}x+b_{i}\right)\right)-\partial_{j}g(x)\right)^{2}{\,\mathrm{d}}x\mathrm{d}\mu^{\otimes k}$
$=\frac{1}{k^{2}}\int_{\Omega}\int_{\left(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right)^{k}}\left(\sum_{i=1}^{k}\left(a_{i}\left\langle w_{i},e_{j}\right\rangle\sigma^{\prime}\left(w_{i}^{\top}x+b_{i}\right)-\partial_{j}g(x)\right)\right)^{2}{\,\mathrm{d}}\mu^{\otimes k}\mathrm{d}x$
$=\frac{1}{k^{2}}\int_{\Omega}\sum_{i=1}^{k}\int_{\left(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right)^{k}}\left(a_{i}\left\langle w_{i},e_{j}\right\rangle\sigma^{\prime}\left(w_{i}^{\top}x+b_{i}\right)-\partial_{j}g(x)\right)^{2}{\,\mathrm{d}}\mu^{\otimes k}\mathrm{d}x$
$=\frac{1}{k}\int_{\Omega}\int_{\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}}\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)-\partial_{j}g(x)\right)^{2}{\,\mathrm{d}}\mu(a,w,b)\mathrm{d}x$
$=\frac{1}{k}\int_{\Omega}\operatorname{Var}_{\mu(a,w,b)}\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)\right){\,\mathrm{d}}x$
$\leq\frac{1}{k}\int_{\Omega}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)\right)^{2}\right]{\,\mathrm{d}}x.$
Since $\sigma$ is the cosine function, we have
$\mathbb{E}_{\mu^{\otimes k}}\|\nabla R(\Theta;x)\|_{L^{2}(\Omega)}^{2}$ $=\mathbb{E}_{\mu^{\otimes k}}\sum_{j=1}^{d}\left\|\partial_{j}R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$
$\leq\frac{1}{k}\int_{\Omega}\sum_{j=1}^{d}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)\right)^{2}\right]{\,\mathrm{d}}x$
$\leq\frac{1}{k}\int_{\Omega}\sum_{j=1}^{d}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\right)^{2}\right]{\,\mathrm{d}}x=\frac{1}{k}\int_{\Omega}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right]{\,\mathrm{d}}x$
$\leq\frac{m(\Omega)}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right]=\frac{1}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right].$ (3.37)
Combining (3.33), (3.36), and (3.3), we obtain
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{H^{1}(\Omega)}$ $\leq\frac{1}{k}\left(\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\right]+\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\|w\|^{2}\right]\right)$
$=\frac{1}{k}\sum_{\omega\in\mathbb{Z}^{d}}\frac{|c_{\omega}|}{Z}Z^{2}\left(1+\pi^{2}\|\omega\|^{2}\right)=\frac{Z}{k}\left(\sum_{\omega\in\mathbb{Z}^{d}}|c_{\omega}|\left(1+\pi^{2}\|\omega\|^{2}\right)\right)$
$\leq\frac{\left(\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}+\varepsilon\right)^{2}}{k}.$
Letting $\varepsilon\to 0$ yields the desired estimate (3.32), which completes the proof of the theorem. ∎
Proof of Theorem 2.13.
The conclusion follows by combining Proposition 2.10, Corollary 3.7, Theorem 3.9, Proposition 3.11, and Theorem 3.14. ∎
In order to prove the approximation result using the $\operatorname{ReLU}$ activation, we state Theorem 3.15, which is a revised version of in [LLW21, Theorem 17] adapted to the setting of our Barron norm definition. The Barron norm in that work is defined in a way similar to our $c$-Barron norm, but it is based on the $\ell^{1}$-norm $\|k\|_{1}$ rather than the Euclidean norm $\|k\|$ used here. Consequently, their Barron space is substantially smaller than ours. Nevertheless, Lemma 18 and Proposition 19 in [LLW21]—which are key ingredients in the proof of Theorem 17 there—can be extended to our setting by adapting their arguments with suitable modifications. We provide the complete proof in Appendix A.4 for completeness.
Theorem 3.15.
Suppose $g\in{\mathcal{B}}_{s}^{2}(\Omega)$ is an $s$-Barron function of weight 2. Then, for any fixed positive integer $k$, there exist a constant $c\in\mathbb{R}$ with $|c|\leq 2\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}$, and a collection of parameters
$\left\{(a_{i},w_{i},b_{i})\in\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right\}_{1\leq i\leq k}$
satisfying
[…]
such that
$\left\|c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i})-g(x)\right\|_{H^{1}(\Omega)}\leq\frac{\sqrt{q(d)}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}}{\sqrt{k}},$
where $q(d)=256d^{2}+128d+4$. An analogous result also holds if ${\mathcal{B}}_{s}^{2}(\Omega)$ is replaced by ${\mathcal{B}}_{c}^{2}(\Omega)$.
Proof of Theorem 2.17.
The result follows by combining Corollary 3.7, Theorem 3.9, Proposition 3.11, and Theorem 3.15. ∎
Appendix A Omitted Proofs in Sections 2 and 3
A.1 Proof of Theorem 2.4
We denote the domain $(-1,1)^{d}\subseteq\mathbb{R}^{d}$ by $\widetilde{\Omega}$.
[…]
$\left\{\prod_{i=1}^{d}\operatorname{sc}(\pi k_{i}x_{i}):k=(k_{1},\ldots,k_{d})\in\mathbb{N}^{d},\ x=(x_{1},\ldots,x_{d})\in\widetilde{\Omega}\right\}$, (A.1)
[…]
$\widetilde{g}(x)\coloneqq\begin{cases}\operatorname{sgn}(x_{1})\cdots\operatorname{sgn}(x_{d})g(|x_{1}|,\dots,|x_{d}|),&\text{if }x_{m}\neq 0\text{ for all }m=1,\dots,d,\\ 0,&\text{otherwise},\end{cases}$
where $\operatorname{sgn}(t)$ denotes the sign function, defined by
$\operatorname{sgn}(t)=\begin{cases}1,&\text{if }t>0,\\ 0,&\text{if }t=0,\\ -1,&\text{if }t<0.\end{cases}$
By construction, $\widetilde{g}$ is odd in each coordinate, and hence its expansion in the basis (A.1) contains only sine terms:
[…]
Restricting this expansion to $\Omega$ yields a representation of $g$ in terms of the collection in (2). The uniqueness of the expansion is ensured by the orthogonality of distinct functions in this system, and the coefficients are real since
$a_{k}=2^{d}\int_{\Omega}g(x)S_{k}(x)\,\mathrm{d}x\in\mathbb{R}.$
[…]
$\widetilde{g}(x)\coloneqq\begin{cases}g(|x_{1}|,\dots,|x_{d}|),&\text{if }x_{m}\neq 0\text{ for all }m=1,\dots,d,\\ 0,&\text{otherwise}.\end{cases}$
By construction, $\widetilde{g}$ is even in each coordinate, and hence its expansion in the basis (A.1) contains only cosine terms:
[…]
The remaining part of the proof follows the same argument as in (2).
[…]
$\widetilde{g}(x)\coloneqq\begin{cases}\operatorname{sgn}(x_{i})\operatorname{sgn}(x_{j})g(|x_{1}|,\dots,|x_{d}|),&\text{if }x_{m}\neq 0\text{ for all }m=1,\dots,d,\\ 0,&\text{otherwise}.\end{cases}$
By construction, $\widetilde{g}$ is odd in the $i$-th and $j$-th coordinates and even in the remaining ones. As a result, its expansion in the basis (A.1) consists of sine terms in the $i$-th and $j$-th positions and cosine terms in the remaining positions:
[…]
The remaining part of the proof follows the same argument as in (2).
Example A.1.
We present an example demonstrating that a function in $L^{2}(\Omega)$ may not have a unique expansion in terms of the exponential basis (1).
Consider the constant function $g(x)=1$ on $\Omega$. Then $g$ admits at least two distinct expansions with respect to the exponential basis:
$g(x)=e^{i\pi 0^{\top}x}=\sum_{\omega\in\mathbb{Z}^{d}}a_{\omega}e^{i\pi\omega^{\top}x}\quad\text{in }L^{2}(\Omega),$
where one possible choice of the coefficients $\{a_{\omega}\}$ is given by
$a_{\omega}=\begin{cases}\displaystyle\frac{1}{2^{d-1}}-1,&\omega=0,\\ \displaystyle\frac{1}{2^{d-1}}\prod_{1\leq j\leq d,\,\omega_{j}\neq 0}\frac{1-e^{-i\pi\omega_{j}}}{i\pi\omega_{j}},&\text{otherwise}.\end{cases}$
[…]
A.2 Proof of Lemma 3.6
Let $A(x)=O(x)\Lambda(x)O(x)^{\top}$ be the orthogonal decomposition of $A(x)$, where $O(x)$ is an orthogonal matrix and $\Lambda(x)=\mathrm{diag}\{\lambda_{1}(x),\dots,\lambda_{d}(x)\}$ is a diagonal matrix whose diagonal entries are the eigenvalues of $A(x)$. We claim that
$a_{\max}=\sup_{x\in\Omega}\sup_{1\leq i\leq d}|\lambda_{i}(x)|.$
Indeed, since orthogonal transformations preserve the Euclidean norm, we have
$a_{\max}$ $=\|A(x)\|_{\mathrm{op}}=\sup_{x\in\Omega}\sup_{v\in\mathbb{R}^{d}\setminus\{0\}}\frac{\|A(x)v\|}{\|v\|}=\sup_{x\in\Omega}\sup_{v\in\mathbb{R}^{d}\setminus\{0\}}\frac{\|O(x)\Lambda(x)O(x)^{\top}v\|}{\|O(x)^{\top}v\|}$
$=\sup_{x\in\Omega}\sup_{\xi\in\mathbb{R}^{d}\setminus\{0\}}\frac{\|\Lambda(x)\xi\|}{\|\xi\|}=\sup_{x\in\Omega}\sup_{\xi\in\mathbb{R}^{d}\setminus\{0\}}\left(\frac{\lambda_{1}(x)^{2}\xi_{1}^{2}+\cdots+\lambda_{d}(x)^{2}\xi_{d}^{2}}{\xi_{1}^{2}+\cdots+\xi_{d}^{2}}\right)^{1/2}$
$=\sup_{x\in\Omega}\sup_{1\leq i\leq d}|\lambda_{i}(x)|.$
Therefore, by the Cauchy–Schwarz inequality, we obtain
$\int_{\Omega}\nabla u\cdot A\nabla v{\,\mathrm{d}}x$ $=\int_{\Omega}(\nabla uO(x))\cdot\Lambda(x)\cdot(\nabla vO(x))^{\top}{\,\mathrm{d}}x$
$\leq a_{\max}\|\nabla uO(x)\|_{L^{2}(\Omega)}\|\nabla vO(x)\|_{L^{2}(\Omega)}$
$=a_{\max}\|\nabla u\|_{L^{2}(\Omega)}\|\nabla v\|_{L^{2}(\Omega)}.$
A.3 The Poincaré Constant for the Unit Hypercube
Theorem A.2.
The Poincaré constant $C_{P}$ for the unit hypercube $\Omega$ is $1/\pi\sqrt{d}$; that is,
$C_{P}=\frac{1}{\pi\sqrt{d}}$
is the optimal constant such that
$\|\phi\|_{L^{2}(\Omega)}\leq C_{P}\|\nabla\phi\|_{L^{2}(\Omega)}\quad\text{for all }\phi\in H_{0}^{1}(\Omega).$
Proof.
It is easy to see that
$\frac{1}{C_{P}}=\inf_{\begin{subarray}{c}\phi\in H_{0}^{1}(\Omega)\\ \phi\neq 0\end{subarray}}\frac{\|\nabla\phi\|_{L^{2}(\Omega)}}{\|\phi\|_{L^{2}(\Omega)}}.$
This quantity is the square root of the smallest eigenvalue of the Dirichlet Laplacian (cf. [Eva10, Theorem 6.5.2]):
$\Delta u=\lambda u\quad\text{in }\Omega,\quad u=0\quad\text{on }\partial\Omega.$ (A.2)
It is well known that, for $\Omega=(0,1)^{d}$, the eigenfunctions of the Dirichlet Laplacian are
$\left\{S_{k}(x)=\prod_{i=1}^{d}\sin(\pi k_{i}x_{i}):k=(k_{1},\ldots,k_{d})\in\mathbb{N}^{d},\ k_{i}\neq 0\text{ for all }i,\ x=(x_{1},\ldots,x_{d})\in\Omega\right\},$
and the corresponding eigenvalue for $S_{k}(x)$ is $\pi^{2}\|k\|^{2}$ (cf. [CH89, Chapter 6.4.1]). The smallest eigenvalue of (A.2) is therefore $\pi^{2}d$. Hence,
$C_{P}=\frac{1}{\sqrt{\pi^{2}d}}=\frac{1}{\pi\sqrt{d}}.$
A.4 Proof of Theorem 3.15
As mentioned in Section 3.3, our strategy is a revision and generalization of the methods in [LLW21]. In particular, Lemma A.3 generalizes their Lemma 18, Lemma A.4 generalizes their Proposition 19, and Lemma A.5 is simply a restatement of their Lemma 16.
We let $I_{d}\coloneqq[-\sqrt{d},\sqrt{d}]$ throughout the arguments.
Lemma A.3.
Suppose $g\in C^{2}(I_{d})$ and there exists $B>0$ such that
$\|g^{(s)}\|_{L^{\infty}(I_{d})}\leq B\quad\text{for all }s=0,1,2,$
and there exists $\alpha\in(-\sqrt{d},\sqrt{d})$ such that $g^{\prime}(\alpha)=0$. Then, for any positive integer $m$, there exists a function $g_{m}(z)$ defined on $I_{d}$ of the form
$g_{m}(z)=c+\sum_{i=1}^{2m}a_{i}\operatorname{ReLU}(\varepsilon_{i}z+b_{i}),$ (A.3)
where
$|c|\leq B,\quad|a_{i}|\leq\frac{4\sqrt{d}B}{m},\quad\varepsilon_{i}\in\{\pm 1\},\quad|b_{i}|\leq\sqrt{d},\quad\text{for all }1\leq i\leq k,$ (A.4)
such that
$\|g-g_{m}\|_{H^{1}(I_{d})}\leq\frac{\sqrt{10}d^{5/4}B}{m}.$ (A.5)
Proof.
Let $\{z_{j}\}_{0\leq j\leq 2m}$ be a partition of $I_{d}$ with
$z_{0}=-\sqrt{d},\quad z_{m}=\alpha,\quad z_{2m}=\sqrt{d},$
$z_{j+1}-z_{j}=h_{1}\coloneqq\frac{\alpha+\sqrt{d}}{m},\quad j=0,\ldots,m-1,$
$z_{j+1}-z_{j}=h_{2}\coloneqq\frac{\sqrt{d}-\alpha}{m},\quad j=m,\ldots,2m-1.$
Let $g_{m}(z)$ defined on $I_{d}$ be the piecewise linear interpolation of $g(z)$ with respect to $\{z_{j}\}_{0\leq j\leq 2m}$, i.e.,
$g_{m}(z)\coloneqq\begin{cases}g(z_{j+1})\dfrac{z-z_{j}}{h_{1}}+g(z_{j})\dfrac{z_{j+1}-z}{h_{1}},&z\in[z_{j},z_{j+1}],\ j=0,\ldots,m-1,\\[5.16663pt] g(z_{j+1})\dfrac{z-z_{j}}{h_{2}}+g(z_{j})\dfrac{z_{j+1}-z}{h_{2}},&z\in[z_{j},z_{j+1}],\ j=m,\ldots,2m-1.\end{cases}$
Then, according to [AG11, Chapter 11], and using $h_{1},h_{2}\leq 2\sqrt{d}/m$, we obtain
$\|g-g_{m}\|_{L^{\infty}(I_{d})}\leq\frac{\max\{h_{1},h_{2}\}^{2}}{8}\,\|g^{\prime\prime}\|_{L^{\infty}(I_{d})}\leq\frac{dB}{2m^{2}}\leq\frac{dB}{m}.$ (A.6)
We also claim that
$\|g^{\prime}-g_{m}^{\prime}\|_{L^{\infty}(I_{d})}\leq\frac{2\sqrt{d}B}{m}\leq\frac{2dB}{m}.$ (A.7)
Indeed, if $z\in[z_{j},z_{j+1}]$ for some $0\leq j\leq m-1$, then by the mean value theorem there exist $\xi,\eta\in(z_{j},z_{j+1})$ such that
$|g^{\prime}(z)-g_{m}^{\prime}(z)|=\left|g^{\prime}(z)-\frac{g(z_{j+1})-g(z_{j})}{h_{1}}\right|=|g^{\prime}(z)-g^{\prime}(\xi)|=|g^{\prime\prime}(\eta)||z-\xi|\leq|g^{\prime\prime}(\eta)|h_{1}\leq\frac{2\sqrt{d}B}{m}.$
The same bound follows by the same argument when $z\in[z_{j},z_{j+1}]$ for some $m\leq j\leq 2m-1$. This proves (A.7). The estimate (A.5) then follows by combining (A.6), (A.7), and
$\|g-g_{m}\|_{H^{1}(I_{d})}^{2}=\|g-g_{m}\|_{L^{2}(I_{d})}^{2}+\|g^{\prime}-g_{m}^{\prime}\|_{L^{2}(I_{d})}^{2}\leq 2\sqrt{d}\|g-g_{m}\|_{L^{\infty}(I_{d})}^{2}+2\sqrt{d}\|g^{\prime}-g_{m}^{\prime}\|_{L^{\infty}(I_{d})}^{2}.$
Now we show that $g_{m}(z)$ can be written in the ReLU form (A.3) and satisfies the coefficient bound in (A.4). Indeed, it is straightforward to verify that
$g_{m}(z)=g(z_{m})+\sum_{i=1}^{m}a_{i}\operatorname{ReLU}(-z+z_{i})+\sum_{i=m+1}^{2m}a_{i}\operatorname{ReLU}(z-z_{i-1}),\quad z\in I_{d},$
where
$a_{i}=\begin{cases}\dfrac{g(z_{i-1})-2g(z_{i})+g(z_{i+1})}{h_{1}},&i=1,\ldots,m-1,\\[4.30554pt] \dfrac{g(z_{m-1})-g(z_{m})}{h_{1}},&i=m,\\[4.30554pt] \dfrac{g(z_{m+1})-g(z_{m})}{h_{2}},&i=m+1,\\[4.30554pt] \dfrac{g(z_{i-2})-2g(z_{i-1})+g(z_{i})}{h_{2}},&i=m+2,\ldots,2m.\end{cases}$
It remains to verify that $|a_{i}|$ satisfies the upper bound in (A.4).
Case $i=1,\ldots,m-1$. By the Taylor expansion, there exist $\eta_{1}\in(z_{i-1},z_{i})$ and $\eta_{2}\in(z_{i},z_{i+1})$ such that
$g(z_{i-1})$ $=g(z_{i})-g^{\prime}(z_{i})h_{1}+g^{\prime\prime}(\eta_{1})h_{1}^{2},$
$g(z_{i+1})$ $=g(z_{i})+g^{\prime}(z_{i})h_{1}+g^{\prime\prime}(\eta_{2})h_{1}^{2}.$
Hence,
$|a_{i}|=\left|\frac{g(z_{i-1})-2g(z_{i})+g(z_{i+1})}{h_{1}}\right|=|g^{\prime\prime}(\eta_{1})h_{1}+g^{\prime\prime}(\eta_{2})h_{1}|\leq 2Bh_{1}\leq\frac{4\sqrt{d}B}{m}.$
Case $i=m$. By the construction of the partition, $g^{\prime}(z_{m})=g^{\prime}(\alpha)=0$. Hence, by the mean value theorem there exist $\xi,\eta\in(z_{m-1},z_{m})$ such that
$|a_{m}|=\left|\frac{g(z_{m-1})-g(z_{m})}{h_{1}}\right|=|g^{\prime}(\xi)|=|g^{\prime}(\xi)-g^{\prime}(z_{m})|=|g^{\prime\prime}(\eta)|\,h_{1}\leq\frac{2\sqrt{d}B}{m}.$
Case $i=m+1$. By a similar argument to that in the case $i=m$, we obtain the same bound.
Case $i=m+2,\ldots,2m$. By a similar argument to that in the case $1\leq i\leq m-1$, we obtain the same bound. ∎
Lemma A.4.
For $B>0$, define the following families of functions on $\Omega=(0,1)^{d}\subseteq{\mathbb{R}}^{d}$ (it is easy to verify that they are all in $H^{1}(\Omega)$):
$\mathscr{F}_{\sin}(B)$ $\coloneqq\left\{\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi k^{\top}x):|A|\leq B,\ k\in\mathbb{N}^{d}\setminus\{0\}\right\},$
$\mathscr{F}_{\cos}(B)$ $\coloneqq\left\{\frac{A}{1+\pi^{2}\|k\|^{2}}\cos(\pi k^{\top}x):|A|\leq B,\ k\in\mathbb{N}^{d}\setminus\{0\}\right\},$
$\mathscr{F}_{\operatorname{ReLU}}(B)$ $\coloneqq\left\{c+A\operatorname{ReLU}(w^{\top}x+b):|c|\leq B,\ |A|\leq 8\sqrt{d}B,\ \|w\|=1,\ |b|\leq\sqrt{d}\right\}.$
Then the $H^{1}(\Omega)$-closures of the convex hulls of $\mathscr{F}_{\sin}(B)$ and $\mathscr{F}_{\cos}(B)$ are both contained in the $H^{1}(\Omega)$-closure of the convex hull of $\mathscr{F}_{\operatorname{ReLU}}(B)$.
Proof.
We only prove the sine function case, as the cosine case follows similarly. It suffices to show that
$\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi k^{\top}x)\in\text{the $H^{1}(\Omega)$-closure of }\operatorname{conv}\left(\mathscr{F}_{\operatorname{ReLU}}(B)\right).$
By Lemma A.3, the function $\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi\|k\|z)$ defined on $I_{d}$ can be $H^{1}(I_{d})$-approximated by a linear combination of a constant and terms of the form $\operatorname{ReLU}(\varepsilon z+b)$, with the sum of the absolute values of the coefficients of $\operatorname{ReLU}(\varepsilon z+b)$ bounded by $8\sqrt{d}B$. Hence, $\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi\|k\|z)$ is contained in the $H^{1}(I_{d})$-closure of
$\operatorname{conv}\left\{c+A\operatorname{ReLU}(\varepsilon z+b):|c|\leq B,\ |A|\leq 8\sqrt{d}B,\ \varepsilon\in\{\pm 1\},\ |b|\leq\sqrt{d}\right\}.$
Since $\|(k/\|k\|)^{\top}x\|\leq\|x\|\leq\sqrt{d}$, it follows that
$\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi k^{\top}x)=\frac{A}{1+\pi^{2}\|k\|^{2}}\sin\!\left(\pi\|k\|\,(k/\|k\|)^{\top}x\right)$
also lies in the $H^{1}(\Omega)$-closure of
$\operatorname{conv}\left\{c+A\operatorname{ReLU}\left(\varepsilon(k/\|k\|)^{\top}x+b\right):|c|\leq B,\ |A|\leq 8\sqrt{d}B,\ \varepsilon\in\{\pm 1\},\ |b|\leq\sqrt{d}\right\}.$ (A.8)
The result then follows from the fact that (A.8) is contained in $\operatorname{conv}\left(\mathscr{F}_{\operatorname{ReLU}}(B)\right)$. ∎
Lemma A.5 ([Pis81, Bar93]).
Let $\mathscr{G}$ be a subset of a Hilbert space such that the norm of every element in $\mathscr{G}$ is bounded by $B_{\mathscr{G}}>0$. Suppose $u$ belongs to the closure of the convex hull of $\mathscr{G}$. Then, for every positive integer $k$, there exist $\{g_{i}\}_{i=1}^{k}\subseteq\mathscr{G}$ and $\{c_{i}\}_{i=1}^{k}\subseteq[0,1]$ with $\sum_{i=1}^{k}c_{i}=1$ such that
$\left\|u-\sum_{i=1}^{k}c_{i}g_{i}\right\|\leq\frac{B_{\mathscr{G}}}{\sqrt{k}}.$
Proof of Theorem 3.15.
We only prove the case for ${\mathcal{B}}_{s}^{2}(\Omega)$, as the case for ${\mathcal{B}}_{c}^{2}(\Omega)$ follows similarly.
Although the definition of an $s$-Barron function involves the basis $\{S_{k}(x)\}$ with $k=(k_{1},\ldots,k_{d})\in\mathbb{N}^{d}$ such that all $k_{i}\neq 0$, for the purpose of unifying the proof with the ${\mathcal{B}}_{c}^{2}(\Omega)$ case, we set $a_{k}=0$ for all $k\in\mathbb{N}^{d}$ having at least one component $k_{i}=0$ in the expansion
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x).$
With this convention, we can rewrite
$g(x)-a_{0}=\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}a_{k}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}a_{k}\cdot\frac{1}{2^{d}}\sum_{\widetilde{k}\in\{(\pm k_{1},\ldots,\pm k_{d})\}}\varepsilon\operatorname{sc}(\pi\widetilde{k}^{\top}x)$
$=$ $\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}\sum_{\widetilde{k}\in\{(\pm k_{1},\ldots,\pm k_{d})\}}\frac{\varepsilon a_{k}\,(1+\pi^{2}\|\widetilde{k}\|^{2})}{2^{d}A_{g}}\cdot\frac{A_{g}}{1+\pi^{2}\|\widetilde{k}\|^{2}}\,\operatorname{sc}(\pi\widetilde{k}^{\top}x).$
Here $\varepsilon\in\{\pm 1\}$, $\operatorname{sc}$ denotes $\cos$ if $d$ is even and $\sin$ if $d$ is odd, and
$A_{g}\coloneqq\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}|a_{k}|\,\left(1+\pi^{2}\|k\|^{2}\right)\leq\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}.$
(In the ${\mathcal{B}}_{c}^{2}(\Omega)$ case, only the cosine term appears.) Under this normalization, we have
$\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}\sum_{\widetilde{k}\in\{(\pm k_{1},\ldots,\pm k_{d})\}}\left|\frac{\varepsilon a_{k}\,(1+\pi^{2}\|\widetilde{k}\|^{2})}{2^{d}A_{g}}\right|=1,$
Next, we observe that
$\|g\|_{H^{1}(\Omega)}^{2}=\sum_{k\in\mathbb{N}^{d}}|\alpha_{k}|(1+\pi^{2}\|k\|^{2})|a_{k}|^{2}\leq\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}^{2}<\infty,$
where $\alpha_{k}$ is a constant depending on $k$ with $|\alpha_{k}|\leq 1$. This shows that $g(x)\in H^{1}(\Omega)$, and hence $g(x)-a_{0}\in H^{1}(\Omega)$ as well.
Combining the above observations, we conclude that $g(x)-a_{0}$ lies in the $H^{1}(\Omega)$-closure of either $\operatorname{conv}(\mathscr{F}_{\cos}(\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}))$ or $\operatorname{conv}(\mathscr{F}_{\sin}(\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}))$. Hence, by Lemma A.4, it is contained in the $H^{1}(\Omega)$-closure of $\operatorname{conv}(\mathscr{F}_{\operatorname{ReLU}}(\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}))$. Moreover, since $|a_{0}|\leq\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}$ (not needed here as $a_{0}=0$, but required for the ${\mathcal{B}}_{c}^{2}(\Omega)$ case; we retain it for consistency of the proofs), it follows that $g(x)$ lies in the $H^{1}(\Omega)$-closure of the convex hull of
$\mathscr{G}\coloneqq\left\{c+A\operatorname{ReLU}(w^{\top}x+b):|c|\leq 2\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)},\ |A|\leq 8\sqrt{d}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)},\ \|w\|=1,\ |b|\leq\sqrt{d}\right\}.$
The result then follows from Lemma A.5 together with the bound for the $H^{1}$-norm of each $h\in\mathscr{G}$:
$\|h\|_{H^{1}(\Omega)}^{2}$ $\leq\left(2\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}+8\sqrt{d}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}(\sqrt{d}+\sqrt{d})\right)^{2}+(8\sqrt{d}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)})^{2}$
$=(256d^{2}+128d+4)\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}^{2}.$
Acknowledgements
The work of ZC is supported in part by the National Science Foundation via grant DMS-2509011. ZC thanks Jianfeng Lu and Yulong Lu for helpful discussions. LH gratefully acknowledges Padmavathi Srinivasan for her continuous encouragement and patient guidance during his first two years of graduate study. He also thanks Wenkui Du and Shengxuan Zhou for many valuable discussions.