### v2

# Barron Space Representations for Elliptic PDEs with Homogeneous Boundary Conditions

Abstract
We study the approximation complexity of high-dimensional second-order elliptic PDEs with homogeneous boundary conditions on the unit hypercube, within the framework of Barron spaces. Under the assumption that the coefficients belong to suitably defined Barron spaces, we prove that the solution can be efficiently approximated by two-layer neural networks, circumventing the curse of dimensionality. Our results demonstrate the expressive power of shallow networks in capturing high-dimensional PDE solutions under appropriate structural assumptions.
Contents
1 Introduction
High-dimensional partial differential equations (PDEs) arise in a wide range of applications, including physics, finance, and control. Classical numerical schemes, such as finite difference and finite element methods, become computationally intractable as the spatial dimension $d$ increases, due to the so-called curse of dimensionality (CoD): to achieve an accuracy of $\varepsilon$, the computational cost scales as $\mathcal{O}(\varepsilon^{-d})$, exponentially in $d$.
To overcome this barrier, a variety of deep learning-based approaches have been developed in recent years [EHJ17, HJE18, SS18, EY18, HZE19, RPK19, HSN20, PSMF20, ZHL21, NZGK21, EHJ22, KT24, KT25]. As an example, the Deep Ritz method (DRM) proposed by E and Yu [EY18] is a powerful approach for solving PDEs that admit a variational formulation. In this method, the PDE solution is represented by a deep neural network, and the network parameters are determined by minimizing the associated energy functional, with the integrals evaluated via Monte Carlo sampling. Since only first-order derivatives of the network output are required, this approach is particularly efficient for PDEs involving higher-order derivatives. Additionally, the Physics-Informed Neural Networks (PINNs) developed by Raissi et al. [RPK19] embeds the PDE and boundary conditions into the loss function by penalizing violations at selected collocation points. While this framework applies to a broader class of PDEs, it typically requires computing higher-order derivatives of the network output, which can reduce efficiency for PDEs with high-order terms.
Despite differences in formulation, both DRM and PINNs share the goal of learning PDE solutions with neural networks. Their practical performance hinges on several theoretical factors: the representational capacity of the network architecture, the design and convergence of the training algorithm, and the generalization ability of the trained model to unseen inputs [SDK20, LLW21, LL22, DRM24]. These considerations motivate three fundamental theoretical questions for deep learning–based PDE methods:
  • Approximation: What is the minimal architectural complexity (e.g., depth, width) required to approximate a target function to a prescribed accuracy?
  • Optimization: For a given architecture, how can we design training algorithms with provable convergence to an optimal or near-optimal solution?
  • Generalization: Under what conditions does the trained network generalize effectively, achieving low error on unseen data?
In this work, we focus on the first question. A prominent framework for addressing this problem is based on Barron functions. In his seminal work [Bar93], Barron proved that for any real-valued function $f\colon\mathbb{R}^{d}\to\mathbb{R}$, if its Fourier transform $\widehat{f}$ satisfies
$\int_{\mathbb{R}^{d}}\|\xi\|\cdot|\widehat{f}(\xi)|{\,\mathrm{d}}\xi<\infty,$ (1.1)
a condition now referred to as having finite Barron norm, then $f$ can be approximated in the $L^{2}$-norm by a two-layer neural network of the form
$\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma(w_{i}^{\top}x+b_{i}),$ (1.2)
with an approximation error independent of the input dimension $d$. This result provides a rigorous theoretical foundation for the ability of neural networks to circumvent the CoD in high-dimensional function approximation; that is, the required number of neurons does not grow exponentially with the input dimension. Following [Bar93], there have been multiple variants and generalizations of Barron functions with dimension-independent approximation results of a similar spirit, see e.g. [KB18, EMW22, EW22a, SX22, SX23, SX24].
This observation inspires a new paradigm for the study of high-dimensional PDEs:
If the coefficients of a PDE are such that its solution can be well approximated by a Barron function without incurring CoD, then a two-layer neural network can be used to efficiently approximate the solution in high dimensions.
This perspective was first realized by Chen et al. [CLL21], who developed a Barron function-based approximation framework for second-order elliptic PDEs on the whole space $\mathbb{R}^{d}$. Their iterative schemes, motivated by [MLR21], converge exponentially fast while maintaining control of the Barron norm. On bounded domains, Marwah et al. [MLLR23] extended this methodology to Dirichlet problems on the unit hypercube, though their results require additional regularity assumptions on the coefficients (see the discussion of Theorem 2.7 in Section 2.2). In contrast, the case of second-order elliptic PDEs with Neumann boundary conditions on bounded domains remains largely open.
1.1 Our Contributions
We study the approximation of solutions to second-order elliptic PDEs on the $d$-dimensional unit hypercube $\Omega\coloneqq(0,1)^{d}\subseteq\mathbb{R}^{d}$. Specifically, we consider the following two prototypical elliptic problems with homogeneous boundary conditions:
  • The Dirichlet Problem:
[…]
  • The Neumann Problem:
[…]
where $A(x)$ is a real matrix-valued function, and $\nu$ denotes the outward unit normal vector on $\partial\Omega$. Our main results are stated informally below; precise formulations are given in Theorems 2.7 and 2.9.
Theorem 1.1 (Main Results, Informal Version).
Suppose the coefficients $A(x)$, $c(x)$, and $f(x)$ in equations (1.3) and (1.4) are suitable Barron functions. Then the weak solution $u^{*}$ to each equation can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network of the form
$\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i})\quad\text{or}\quad c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i}),$
using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where $C>0$ depends only on the Barron norms of the PDE coefficients.
We briefly outline the strategy below, with full details and precise notation provided in the subsequent section.
  1. Step 1.
    For each of the equations (1.3) and (1.4), we construct the corresponding Sobolev gradient flow:
    $u_{t+1}=u_{t}-\alpha D\mathcal{E}(u_{t}),$
    where $\mathcal{E}$ is the energy functional associated with (1.3) or (1.4). We show that this scheme converges exponentially fast to the solution $u^{*}$ in the $H^{1}(\Omega)$-norm:
    $\|u_{t}-u^{*}\|_{H^{1}(\Omega)}\lesssim\beta^{t},\quad\text{for all }t\geq 0,$
    for some $0<\beta<1$ depending only on the PDE coefficients (see Equations (3.17) and (3.18)). In fact, the iterative schemes previously proposed for variational PDEs in [CLL21, MLLR23] can also be interpreted as special cases of the Sobolev gradient flow.
  2. Step 2.
    We derive a recursive estimate for the Barron norm of $u_{t+1}$ in terms of that of $u_{t}$, showing that it grows at most at rate $\mathcal{O}(d^{2})$ (see Theorem 3.5). This ensures that all iterates $u_{t}$ remain Barron functions.
  3. Step 3.
    Applying the Barron function approximation theory developed in [EMW22] (see Theorem 3.10), any Barron function $g$ can be approximated by a two-layer neural network with cosine activation, and the $H^{1}(\Omega)$-error is bounded by
    $\left\|\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i})-g\right\|_{H^{1}(\Omega)}\leq\frac{\|g\|_{\mathcal{B}(\Omega)}}{\sqrt{k}},$
    where $\|g\|_{\mathcal{B}(\Omega)}$ is the Barron norm of $g$. For the $\operatorname{ReLU}$ activation case, we generalize the strategies in [LLW21] (see Theorem 3.11) and obtain
    $\left\|c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i})-g\right\|_{H^{1}(\Omega)}\lesssim\frac{\|g\|_{{\mathcal{B}}(\Omega)}}{\sqrt{k}}.$
    These results imply that the weak solution can be efficiently approximated by a two-layer neural network with either cosine or $\operatorname{ReLU}$ activation, without suffering from the CoD.
1.2 Related Works
In the Barron function-based framework, beyond approximating the solution by a Barron function, one may also pose the following question:
If the coefficients of the PDE are sufficiently regular in the Barron sense, does the solution belong to the Barron class?
This viewpoint is often referred to as the regularity problem. It was first explored in [LLW21, LL22, EW22b], where the authors investigated the Poisson equation and the Schrödinger equation on the unit hypercube with Neumann boundary conditions. Subsequently, Chen et al. [CLLZ23] implemented this perspective for the stationary Schrödinger equation on $\mathbb{R}^{d}$. Regularity results for the Hamilton–Jacobi–Bellman equation in the whole space were established in [FL25], and for the electronic Schrödinger equation in [Yse25].
Apart from the Barron function-based approach, another influential deep learning methodology was proposed by E et al. [EHJ17, HJE18], who reformulated certain parabolic PDEs as backward stochastic differential equations (BSDEs), and further interpreted these BSDEs as stochastic control problems. These problems, in turn, can be viewed as instances of model-based reinforcement learning. This SDE-based framework was later extended by Sirignano and Spiliopoulos [SS18]. We refer to [TTY22, Rai23, KT24, LVR24] for additional applications of SDE-based methods to high-dimensional PDEs.
Alternative deep learning-based approaches have also been developed, typically relying on strong structural assumptions on the coefficients of the underlying PDEs. For instance, [GH21, GHJvW23] investigate classes of parabolic and Poisson equations that admit stochastic representations, such as those derived from the Feynman–Kac formula. In contrast, [MLR21] proposes neural network approximation conditions on the coefficients of second-order elliptic PDEs.
1.3 Organization
The rest of this paper is organized as follows. Section 2 present our main results after introducing the Barron spaces and basic properties. The proofs, based on a three-step strategy outlined in Section 1.1, are detailed in Section 3.
2 Main Results
In this section, we present our main results. To ensure the existence and uniqueness of weak solutions to PDEs (1.3) and (1.4), we impose the following minimal assumption on their coefficients. All functions in this paper are real-valued unless otherwise noted.
Assumption 2.1 (Coefficients of Elliptic PDEs).
  1. (1)
    The matrix-valued function $A(x)=(A_{ij}(x))_{1\leq i,j\leq d}$ defined on $\Omega$ is symmetric and uniformly elliptic. By uniform ellipticity, we mean that there exists a constant $a_{\min}>0$ such that
    $v^{\top}A(x)v\geq a_{\min}\|v\|^{2}\quad\text{for all }x\in\Omega\text{ and }v\in\mathbb{R}^{d},$
    where $\|v\|$ denotes the standard Euclidean norm in $\mathbb{R}^{d}$. Moreover, the operator norm of $A(x)$, defined as
    $\|A(x)\|_{\mathrm{op}}\coloneqq\sup_{x\in\Omega}\sup_{v\in\mathbb{R}^{d}\setminus\{0\}}\frac{\|A(x)v\|}{\|v\|},$
    is finite, and we denote $a_{\max}\coloneqq\|A(x)\|_{\mathrm{op}}$.
  2. (2)
    The scalar coefficient function $c(x)$ defined on $\Omega$ satisfies
    $0<c_{\min}\leq c(x)\leq c_{\max}<\infty\quad\text{for all }x\in\Omega,$
  3. (3)
    The source term $f(x)$ belongs to $L^{2}(\Omega)$.
We refer to [Eva10, Lie13] for a comprehensive treatment of weak solutions in Sobolev spaces, and provide a brief summary here for completeness. A function $u_{D}^{*}\in H_{0}^{1}(\Omega)$ is said to be a weak solution of the Dirichlet problem (1.3) if it satisfies
$\int_{\Omega}\left(\nabla u_{D}^{*}\cdot A\nabla v+cu_{D}^{*}v-fv\right){\,\mathrm{d}}x=0\quad\text{for all }v\in H_{0}^{1}(\Omega),$ (2.1)
A function $u_{N}^{*}\in H^{1}(\Omega)$ is said to be a weak solution of the Neumann problem (1.4) if it satisfies
$\int_{\Omega}\left(\nabla u_{N}^{*}\cdot A\nabla v+cu_{N}^{*}v-fv\right){\,\mathrm{d}}x=0\quad\text{for all }v\in H^{1}(\Omega).$ (2.2)
Assumption 2.1 and the Lax–Milgram theorem [Eva10, Theorem 6.2.1] ensure the existence and uniqueness of weak solutions $u_{D}^{*}\in H_{0}^{1}(\Omega)$ and $u_{N}^{*}\in H^{1}(\Omega)$ to (1.3) and (1.4), respectively.
2.1 Barron Spaces
To define the four types of Barron functions considered in this paper, we begin by introducing in Theorem 2.2 four families of functions in $L^{2}(\Omega)$, which serve as building blocks for representing elements in this space. Some of these families form bases of $L^{2}(\Omega)$. The justification is deferred to the Appendix A.1.
Theorem 2.2.
Let $\mathbb{N}\coloneqq\{0,1,2,\dots\}$ denote the set of natural numbers. For $k=(k_{1},\dots,k_{d})\in\mathbb{N}^{d}$, $\omega\in\mathbb{Z}^{d}$, and $x=(x_{1},\dots,x_{d})\in\Omega$, we define the following four families of functions in $L^{2}(\Omega)$:
  1. (1)
    $\left\{e^{i\pi\omega^{\top}x}:\omega\in\mathbb{Z}^{d}\right\}$,
  2. (2)
    $\left\{S_{k}(x)=\prod_{i=1}^{d}\sin(\pi k_{i}x_{i}):k\in\mathbb{N}^{d},\ k_{i}\neq 0\text{ for all }i\right\}$,
  3. (3)
    $\left\{C_{k}(x)=\prod_{i=1}^{d}\cos(\pi k_{i}x_{i}):k\in\mathbb{N}^{d}\right\}$,
  4. (4)
    $\left\{M_{k,i,j}(x)=C_{k}(x)\cdot\frac{\sin(\pi k_{i}x_{i})\sin(\pi k_{j}x_{j})}{\cos(\pi k_{i}x_{i})\cos(\pi k_{j}x_{j})}:k\in\mathbb{N}^{d},\ k_{i},k_{j}\neq 0\right\}$, where $1\leq i\neq j\leq d$ are fixed. That is, each function in this family is constructed such that the $i$-th and $j$-th components involve cosine terms, while all remaining components involve sine terms.
Every function in $L^{2}(\Omega)$ admits an $L^{2}$-expansion in family (1) with coefficients in $\mathbb{C}$, though the representation may be non-unique. Families (2)-(4) form orthogonal bases of $L^{2}(\Omega)$ with coefficients in $\mathbb{R}$.
Definition 2.3 (Barron Norms and Barron Functions).
Suppose $g(x)\in L^{2}(\Omega)$.
  1. (1)
    We say that $g$ is an $e$-Barron function of weight $n\geq 0$ if its weighted $e$-Barron norm
    $\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\coloneqq\inf\left\{\sum_{\omega\in\mathbb{Z}^{d}}|a_{\omega}|\left(1+\pi^{n}\|\omega\|^{n}\right):g(x)=\sum_{\omega\in\mathbb{Z}^{d}}a_{\omega}e^{i\pi\omega^{\top}x}\;\text{in }L^{2}(\Omega)\right\}.$
    is finite. The space $\mathcal{B}_{e}^{n}(\Omega)$ is referred to as the $e$-Barron space of weight $n$.
  2. (2)
    We say that $g$ is an $s$-Barron function of weight $n\geq 0$ if it admits the sine basis expansion
    $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)\quad\text{in }L^{2}(\Omega),$
    with finite weighted $s$-Barron norm
    $\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}\coloneqq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right).$
    The space $\mathcal{B}_{s}^{n}(\Omega)$ is referred to as the $s$-Barron space of weight $n$.
  3. (3)
    We say that $g$ is a $c$-Barron function of weight $n\geq 0$ if it admits the cosine basis expansion
    $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}C_{k}(x)\quad\text{in }L^{2}(\Omega),$
    with finite weighted $c$-Barron norm
    $\|g\|_{\mathcal{B}_{c}^{n}(\Omega)}\coloneqq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right).$
    The space $\mathcal{B}_{c}^{n}(\Omega)$ is referred to as the $c$-Barron space of weight $n$.
  4. (4)
    We say that $g$ is an $(i,j)$-mixed Barron function of weight $n\geq 0$ (for fixed indices $i\neq j$) if it admits a mixed basis expansion
    $g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}M_{k,i,j}(x)\quad\text{in }L^{2}(\Omega),$
    with finite weighted $(i,j)$-mixed Barron norm
    $\|g\|_{\mathcal{B}_{i,j}^{n}(\Omega)}\coloneqq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right).$
    The space $\mathcal{B}_{i,j}^{n}(\Omega)$ is referred to as the $(i,j)$-mixed Barron space of weight $n$.
The original definition of the Barron norm for functions on $\mathbb{R}^{d}$ was introduced in [Bar93], as given in equation (1.1). In contemporary literature, the version of the Barron norm defined via the Fourier transform, used in, e.g., [SX22, MLLR23, FL25], is commonly referred to as the spectral Barron norm. In contrast, definitions that avoid the use of the Fourier transform, such as those based on probability measures (e.g., [CLL21, EMW22, SX23]), are typically referred to simply as the Barron norm. Our formulation is inspired by [LLW21, LL22, MLLR23], where cosine basis functions were used in [LLW21, LL22] and exponential basis functions in [MLLR23]. In this work, we introduce the sine-based Barron norm and the $(i,j)$-mixed Barron norm. The sine basis arises naturally when considering functions that vanish on the boundary, while the mixed norm is introduced for certain technical reasons (cf. Lemma 3.8). We also note that if the coefficient matrix $A(x)$ is diagonal, then the mixed Barron norm is not needed. This observation already covers many important families of PDEs, such as the stationary reaction–diffusion equations and static Schrödinger equations.
Note that if $g$ is an $s$-Barron function, then any expansion with respect to the sine basis can be directly rewritten in terms of the exponential basis:
$g(x)$ $=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}}\frac{a_{k}}{(2i)^{d}}\prod_{j=1}^{d}\left(e^{i\pi k_{j}x_{j}}-e^{-i\pi k_{j}x_{j}}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\frac{a_{k}}{(2i)^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\prod_{j=1}^{d}\epsilon_{j}\cdot e^{i\pi(\epsilon\circ k)^{\top}x},$
where $\epsilon\circ k$ is the Hadamard product of $\epsilon$ and $k$. This observation leads to that
$\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\leq\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|\epsilon\circ k\|^{n}\right)=\sum_{k\in\mathbb{N}^{d}}|a_{k}|\left(1+\pi^{n}\|k\|^{n}\right)=\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}.$
A similar argument applies to other types of Barron functions, yielding the following result.
Proposition 2.4.
Let $\widetilde{\mathcal{B}}^{n}(\Omega)$ denote any of the spaces $\mathcal{B}_{s}^{n}(\Omega)$, $\mathcal{B}_{c}^{n}(\Omega)$, or $\mathcal{B}_{i,j}^{n}(\Omega)$. If $g\in\widetilde{\mathcal{B}}^{n}(\Omega)$, then
$\|g\|_{\mathcal{B}_{e}^{n}(\Omega)}\leq\|g\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}.$
In particular, the following inclusions hold:
$\mathcal{B}_{s}^{n}(\Omega),\;\mathcal{B}_{c}^{n}(\Omega),\;\mathcal{B}_{i,j}^{n}(\Omega)\subseteq\mathcal{B}_{e}^{n}(\Omega).$
2.2 Main Results
Our main results are established under the following assumptions on the coefficients appearing in equations (1.3) and (1.4).
Assumption 2.5 (Coefficients of Dirichlet Problem).
  1. (1)
    The entries of the coefficient matrix $A(x)=(A_{ij}(x))_{1\leq i,j\leq d}$ satisfy the following: $A_{ii}(x)\in\mathcal{B}_{c}^{1}(\Omega)$ for all $i$, and $A_{ij}(x)\in\mathcal{B}_{i,j}^{1}(\Omega)$ for all $i\neq j$. We define
    $\ell_{A,D}\coloneqq\max\left\{\max_{1\leq i\leq d}\|A_{ii}\|_{\mathcal{B}_{c}^{1}(\Omega)},\ \max_{1\leq i\neq j\leq d}\|A_{ij}\|_{\mathcal{B}_{i,j}^{1}(\Omega)}\right\}.$
  2. (2)
    The scalar coefficient $c(x)$ lies in $\mathcal{B}_{c}^{2}(\Omega)$, and we denote its Barron norm by $\ell_{c}\coloneqq\|c\|_{\mathcal{B}_{c}^{2}(\Omega)}$.
  3. (3)
    The source function $f(x)$ belongs to $\mathcal{B}_{s}^{0}(\Omega)$, with norm $\ell_{f,D}\coloneqq\|f\|_{\mathcal{B}_{s}^{0}(\Omega)}$.
Assumption 2.6 (Coefficients of Neumann Problem).
  1. (1)
    The entries of the coefficient matrix $A(x)=(A_{ij}(x))_{1\leq i,j\leq d}$ satisfy the following: $A_{ii}(x)\in\mathcal{B}_{s}^{1}(\Omega)$ for all $i$, and $A_{ij}(x)\in\mathcal{B}_{i,j}^{1}(\Omega)$ for all $i\neq j$. We define
    $\ell_{A,N}\coloneqq\max\left\{\max_{1\leq i\leq d}\|A_{ii}\|_{\mathcal{B}_{s}^{1}(\Omega)},\ \max_{1\leq i\neq j\leq d}\|A_{ij}\|_{\mathcal{B}_{i,j}^{1}(\Omega)}\right\}.$
  2. (2)
    The scalar coefficient $c(x)$ lies in $\mathcal{B}_{c}^{2}(\Omega)$, and we denote its Barron norm by $\ell_{c}\coloneqq\|c\|_{\mathcal{B}_{c}^{2}(\Omega)}$.
  3. (3)
    The source function $f(x)$ belongs to $\mathcal{B}_{c}^{0}(\Omega)$, with norm $\ell_{f,N}\coloneqq\|f\|_{\mathcal{B}_{c}^{0}(\Omega)}$.
For consistency in the subsequent analysis, we equip the Hilbert space $H_{0}^{1}(\Omega)$ with the inner product inherited from the standard inner product on $H^{1}(\Omega)$,
$(u,v)_{H^{1}(\Omega)}\coloneqq\int_{\Omega}\nabla u\cdot\nabla v+uv\,\mathrm{d}x.$
Although the norm induced by this inner product differs from the usual one on $H_{0}^{1}(\Omega)$, which is based solely on the gradient term, they are equivalent by Poincaré’s inequality (cf. [Eva10, Theorem 5.6.3]). This choice has no effect on our results.
We now present the first main result of the paper. The precise choice of parameters is given in Remark 2.8, and the full proof is deferred to the next section.
Theorem 2.7 (Main Result 1).
  1. (1)
    Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ denote the weak solution to the Dirichlet problem (1.3). Suppose Assumptions 2.1 and 2.5 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $N_{k}(x)\coloneqq\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i}),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,D}^{2}\left(p_{D}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{D}(d)-1\right)^{2}}\right\rceil,$
    such that $\|N_{k}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
  2. (2)
    Let $u_{N}^{*}\in H^{1}(\Omega)$ denote the weak solution to the Neumann problem (1.4). Suppose Assumptions 2.1 and 2.6 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $N_{k}(x)\coloneqq\frac{1}{k}\sum_{i=1}^{k}a_{i}\cos(w_{i}^{\top}x+b_{i}),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,N}^{2}\left(p_{N}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{N}(d)-1\right)^{2}}\right\rceil,$
    such that $\|N_{k}-u_{N}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
In other words, in both cases, the solution can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network with cosine activation using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where the constant $C>0$ depends only on the Barron norms of PDE coefficients.
Remark 2.8 (Parameter Explanation).
The constants and parameters appearing in Theorem 2.7 are defined as follows. Their derivation and role in the analysis will be discussed in detail in the next section:
$p_{D}(d)=\frac{2+\pi}{2\pi}\alpha_{\ast}\ell_{A,D}d^{2}+\alpha_{\ast}\ell_{c}+1,\qquad p_{N}(d)=\frac{2+\pi}{2\pi}\alpha_{\ast}\ell_{A,N}d^{2}+\alpha_{\ast}\ell_{c}+1,$
$T=\left\lceil\frac{\log\|f\|_{H^{-1}(\Omega)}+|\log\left(\varepsilon\lambda_{\min}/2\right)|}{|\log\beta_{\ast}|}\right\rceil,$
$\alpha_{\ast}=\frac{\lambda_{\min}}{2\lambda_{\max}^{2}},\qquad\beta_{\ast}=\left(1-\frac{\lambda_{\min}^{2}}{4\lambda_{\max}^{2}}\right)^{1/2},$
$\lambda_{\min}=\min\{a_{\min},c_{\min}\},\qquad\lambda_{\max}=\max\{a_{\max},c_{\max}\}.$
The quantities $p_{D}(d)$ and $p_{N}(d)$ can be chosen slightly smaller, although they still scale like $d^{2}$, as suggested by the argument in the proof of Theorem 3.5 in the next section. However, we retain the current definitions to ensure clarity and simplicity of exposition.
Our result is stronger and more general than those in [CLL21, MLLR23] in the following aspects:
  1. (1)
    We do not require the solution to the Dirichlet problem (1.3) to be the restriction of a solution defined on $\mathbb{R}^{d}$, as is needed in [CLL21];
  2. (2)
    The work [MLLR23] claims a comparable network size. However, their derivation relies on the assumption that the coefficients in the Barron expansion vanish except for finitely many terms—an assumption not needed in our analysis;
  3. (3)
    In addition to considering only the Dirichlet problem as in [CLL21, MLLR23], we also provide a result for the Neumann problem.
The effectiveness of periodic activation functions such as $\cos$ and $\sin$ has been empirically demonstrated in recent works [SMB+20, MTS21, BHRZ22, RRG25]. By contrast, for the more commonly used $\operatorname{ReLU}$ activation, analogous approximation results follow from the strategies in [LLW21], and the corresponding proof is also deferred to the next section.
Theorem 2.9 (Main Result 2).
  1. (1)
    Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ denote the weak solution to the Dirichlet problem (1.3). Suppose Assumptions 2.1 and 2.5 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $R_{k}(x)\coloneqq c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i}),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,D}^{2}q(d)\left(p_{D}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{D}(d)-1\right)^{2}}\right\rceil,$
    such that $\|R_{k}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
  2. (2)
    Let $u_{N}^{*}\in H^{1}(\Omega)$ denote the weak solution to the Neumann problem (1.4). Suppose Assumptions 2.1 and 2.6 hold. Then for any $\varepsilon\in(0,2/\lambda_{\min})$, there exists a two-layer neural network of the form
    $R_{k}(x)\coloneqq c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i}),\quad k\leq\left\lceil\frac{\alpha_{\ast}^{2}\ell_{f,N}^{2}q(d)\left(p_{N}(d)^{T}-1\right)^{2}}{\varepsilon^{2}\left(p_{N}(d)-1\right)^{2}}\right\rceil,$
    such that $\|R_{k}-u_{N}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
Here, $q(d)=256d^{2}+128d+4$, and all parameters are as specified in Remark 2.8. Thus, in both cases, the solution can be approximated in the $H^{1}(\Omega)$-norm by a two-layer neural network with $\operatorname{ReLU}$ activation using at most $\mathcal{O}(d^{C|\log\varepsilon|})$ neurons, where the constant $C>0$ depends only on the Barron norms of PDE coefficients.
A similar approximation result for $\operatorname{ReLU}$ activation was also given in [LLW21] with the following differences:
  1. (1)
    Our Barron space is substantially larger than that in [LLW21]. Their Barron norm is defined in a way similar to our $c$-Barron norm, but it is based on the $\ell^{1}$-norm $\|k\|_{1}\coloneqq|k_{1}|+\cdots+|k_{d}|$ rather than the Euclidean norm $\|k\|$ used here;
  2. (2)
    We consider general second-order elliptic PDEs with both Dirichlet and Neumann boundary conditions, rather than only the Poisson equation and the stationary Schrödinger equation with Neumann boundary conditions.
3 Proof of the Main Results
The objective of this section is to prove Theorems 2.7 and 2.9, following the three-step outline presented in Section 1.1. Our argument builds upon the methodology developed in [CLL21, LLW21, MLLR23, CLLZ24].
3.1 Step 1: Sobolev Gradient Flow
3.1.1 The Construction
It is well known that the solutions to the Dirichlet problem (1.3) and the Neumann problem (1.4) in the Sobolev space setting can be characterized as minimizers of suitable energy functionals. This classical characterization is summarized in the following proposition.
Proposition 3.1 (Variational Characterization of Weak Solutions).
  1. (1)
    Suppose Assumption 2.1 holds. Let $u_{D}^{*}\in H_{0}^{1}(\Omega)$ be the weak solution to the Dirichlet problem (1.3). Then $u_{D}^{*}$ is the unique minimizer of the energy functional
    $u_{D}^{*}=\mathop{\operatorname{arg\,min}}_{u\in H_{0}^{1}(\Omega)}\mathcal{E}_{D}(u),\quad\mathcal{E}_{D}(u)\coloneqq\int_{\Omega}\left(\frac{1}{2}\nabla u\cdot A\nabla u+\frac{1}{2}cu^{2}-fu\right)\mathrm{d}x.$ (3.1)
  2. (2)
    Suppose Assumption 2.1 holds. Let $u_{N}^{*}\in H^{1}(\Omega)$ be the weak solution to the Neumann problem (1.4). Then $u_{N}^{*}$ is the unique minimizer of the energy functional
    $u_{N}^{*}=\mathop{\operatorname{arg\,min}}_{u\in H^{1}(\Omega)}\mathcal{E}_{N}(u),\quad\mathcal{E}_{N}(u)\coloneqq\int_{\Omega}\left(\frac{1}{2}\nabla u\cdot A\nabla u+\frac{1}{2}cu^{2}-fu\right)\mathrm{d}x.$ (3.2)
Proof.
We only prove (1), as the proof of (2) follows similarly. By Stampacchia’s theorem (see [Bre11, Theorem 5.6]), it suffices to verify that
$\int_{\Omega}\nabla u_{D}^{*}\cdot A\nabla(v-u_{D}^{*})\,\mathrm{d}x\geq\int_{\Omega}f(v-u_{D}^{*})\,\mathrm{d}x\quad\text{for all }v\in H_{0}^{1}(\Omega).$
This inequality follows directly from the definition of the weak solution given in (2.1). ∎
The classical gradient descent method iteratively updates an initial guess by moving in the direction of the negative gradient of the objective function, aiming to converge to a minimizer. In the context of the PDEs we consider, Proposition 3.1 asserts that the weak solution is the unique minimizer of the associated energy functional. This naturally motivates the application of a gradient descent-type approach to solve the PDE. For the Dirichlet problem, we define the gradient of $\mathcal{E}_{D}$ as the map
$D_{H_{0}^{1}}\mathcal{E}_{D}\colon H_{0}^{1}(\Omega)\longrightarrow H_{0}^{1}(\Omega)$
such that
$\begin{split}\big(D_{H_{0}^{1}}\mathcal{E}_{D}(u),v\big)_{H^{1}(\Omega)}&=\lim_{t\to 0}\frac{\mathcal{E}_{D}(u+tv)-\mathcal{E}_{D}(u)}{t}\\ &=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\end{split}$ (3.3)
for all $u,v\in H_{0}^{1}(\Omega)$. Here we use the inner product on $H_{0}^{1}(\Omega)$ inherited from the standard inner product on $H^{1}(\Omega)$. The uniqueness of $D_{H_{0}^{1}}\mathcal{E}_{D}$ follows from the Riesz representation theorem and we call $D_{H_{0}^{1}}\mathcal{E}_{D}$ the Sobolev gradient of $\mathcal{E}_{D}$ in $H_{0}^{1}(\Omega)$. A similar argument applies to the energy functional $\mathcal{E}_{N}$ for the Neumann problem, and there exists a unique operator $D_{H^{1}}\mathcal{E}_{N}\colon H^{1}(\Omega)\to H^{1}(\Omega)$ such that
$\big(D_{H^{1}}\mathcal{E}_{N}(u),v\big)_{H^{1}(\Omega)}=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }u,v\in H^{1}(\Omega).$ (3.4)
We call $D_{H^{1}}\mathcal{E}_{N}$ the Sobolev gradient of $\mathcal{E}_{N}$ in $H^{1}(\Omega)$. We remark that the concept of the Sobolev gradient has been widely employed in the study of PDEs, including applications to the Gross–Pitaevskii eigenvalue problem [HP20, Zha22, CLLZ24].
Analogous to the fact that the gradient of a function vanishes at its local minimum, the Sobolev gradient of an energy functional should also vanish at its minimizer, i.e.,
$D_{H_{0}^{1}}\mathcal{E}_{D}(u^{*}_{D})=D_{H^{1}}\mathcal{E}_{N}(u^{*}_{N})=0.$ (3.5)
In parallel with the classical gradient descent method, we define iterative schemes based on the Sobolev gradients $D_{H_{0}^{1}}\mathcal{E}_{D}$ and $D_{H^{1}}\mathcal{E}_{N}$ to approximate the solutions of the Dirichlet problem (1.3) and the Neumann problem (1.4), respectively. These schemes are referred to as Sobolev gradient flows, as they can be interpreted as discrete-time approximations of the corresponding continuous Sobolev gradient flows of the energy functionals.
  • Sobolev gradient flow for the Dirichlet problem:
    $u_{t+1}=u_{t}-\alpha D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}),$ (3.6)
    with initial condition $u_{0}=0$, where $\alpha>0$ is a stepsize to be specified later.
  • Sobolev gradient flow for the Neumann problem:
    $u_{t+1}=u_{t}-\alpha D_{H^{1}}\mathcal{E}_{N}(u_{t}),$ (3.7)
    with initial condition $u_{0}=0$, where $\alpha>0$ is a stepsize to be specified later.
3.1.2 Convergence Analysis
This subsection investigates the convergence rate of the Sobolev gradient flows to $u_{D}^{*}$ and $u_{N}^{*}$. In particular, we have the following theorem.
Theorem 3.2.
Assume Assumption 2.1 holds and let $\alpha>0$ be sufficiently small. The Sobolev gradient flow (3.6) for the Dirichlet problem (1.3) converges exponentially in the $H^{1}(\Omega)$-norm to the weak solution $u_{D}^{*}$. More precisely, there exists a constant $0<\beta<1$ such that
$\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\|u_{D}^{*}\|_{H^{1}(\Omega)}\beta^{t}\quad\text{for all }t\geq 0.$
An analogous result holds for the Neumann problem (1.4) under the Sobolev gradient flow (3.7), using the same choice of stepsize $\alpha$ and contraction factor $\beta$.
Proof.
We only present the proof for the Dirichlet problem (1.3), as the proof for the Neumann problem (1.4) follows analogously. Our goal is to establish that there exists a constant $0<\beta<1$ such that
$\|u_{t+1}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\beta\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}\quad\text{for all }t\geq 0.$
By the construction of the Sobolev gradient flow (3.6), we compute
$\begin{split}&\|u_{t+1}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}=\|u_{t}-u_{D}^{*}-\alpha D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}\\ =&\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}+2\alpha\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}+\alpha^{2}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}.\end{split}$ (3.8)
Estimating $(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}))_{H^{1}(\Omega)}$. Let $e_{t}\coloneqq u_{D}^{*}-u_{t}$. By the definition of the energy functional (3.1) and using (3.3), we have
$\mathcal{E}_{D}(u_{D}^{*})-\mathcal{E}_{D}(u_{t})$ $=\mathcal{E}_{D}(e_{t}+u_{t})-\mathcal{E}_{D}(u_{t})$
$=\int_{\Omega}\frac{1}{2}\nabla e_{t}\cdot A\nabla e_{t}+\frac{1}{2}ce_{t}^{2}{\,\mathrm{d}}x+\left(e_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}.$
Hence,
$\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}=\mathcal{E}_{D}(u_{D}^{*})-\mathcal{E}_{D}(u_{t})-\int_{\Omega}\frac{1}{2}\nabla e_{t}\cdot A\nabla e_{t}+\frac{1}{2}ce_{t}^{2}{\,\mathrm{d}}x.$ (3.9)
Since $u_{D}^{*}$ minimizes $\mathcal{E}_{D}$ over $H_{0}^{1}(\Omega)$ by Proposition 3.1, and $u_{t}\in H_{0}^{1}(\Omega)$ for all $t\geq 0$, it follows that $\mathcal{E}_{D}(u_{D}^{*})\leq\mathcal{E}_{D}(u_{t})$. Letting $\lambda_{\min}\coloneqq\min\{a_{\min},c_{\min}\}>0$, we deduce from (3.9) that
$\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}$ $\leq-\int_{\Omega}\frac{1}{2}\nabla e_{t}\cdot A\nabla e_{t}+\frac{1}{2}ce_{t}^{2}{\,\mathrm{d}}x$
$\leq-\frac{1}{2}\lambda_{\min}\int_{\Omega}\|\nabla e_{t}\|^{2}+\|e_{t}\|^{2}{\,\mathrm{d}}x=-\frac{1}{2}\lambda_{\min}\|e_{t}\|_{H^{1}(\Omega)}^{2}.$
Therefore,
$\left(u_{D}^{*}-u_{t},D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}\leq-\frac{1}{2}\lambda_{\min}\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}^{2}.$ (3.10)
Estimating $\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}$. By the vanishing property of the Sobolev gradient at the minimizer (3.5), we have
$\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{D}^{*}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}=0,\quad\text{for all }t\geq 0.$ (3.11)
Let $\lambda_{\max}\coloneqq\max\{a_{\max},c_{\max}\}$. Then, by combining (3.11), the definition of the Sobolev gradient in (3.3), and the Cauchy–Schwarz inequality, we obtain
$\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}^{2}$
$=$ $\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}-\left(D_{H_{0}^{1}}\mathcal{E}_{D}(u_{D}^{*}),D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\right)_{H^{1}(\Omega)}$
$=$ $\int_{\Omega}\nabla(u_{t}-u_{D}^{*})\cdot A\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})+c(u_{t}-u_{D}^{*})D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t}){\,\mathrm{d}}x$
$\leq$ $a_{\max}\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}$
$\qquad\qquad+c_{\max}\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}$
$\leq$ $\lambda_{\max}\Big(\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}$
$\qquad\qquad+\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}\Big)$
$\leq$ $\lambda_{\max}\left(\|\nabla(u_{t}-u_{D}^{*})\|_{L^{2}(\Omega)}^{2}+\|u_{t}-u_{D}^{*}\|_{L^{2}(\Omega)}^{2}\right)^{1/2}$
$\qquad\qquad\cdot\left(\|\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}^{2}+\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{L^{2}(\Omega)}^{2}\right)^{1/2}$
$=$ $\lambda_{\max}\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}.$
This gives
$\|D_{H_{0}^{1}}\mathcal{E}_{D}(u_{t})\|_{H^{1}(\Omega)}\leq\lambda_{\max}\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}.$ (3.12)
Combining the estimates (3.8), (3.10), and (3.12), we obtain
$\|u_{t+1}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\left(1-\lambda_{\min}\alpha+\lambda_{\max}^{2}\alpha^{2}\right)^{1/2}\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}.$
Define the contraction factor
$\beta\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\left(1-\lambda_{\min}\alpha+\lambda_{\max}^{2}\alpha^{2}\right)^{1/2}.$ (3.13)
Then for sufficiently small $\alpha>0$, we can ensure that $0<\beta<1$, which completes the proof. ∎
Let
$\lambda_{\min}\coloneqq\min\{a_{\min},c_{\min}\},\quad\lambda_{\max}\coloneqq\max\{a_{\max},c_{\max}\},$
and define the stepsize and contraction factor by
$\alpha_{\ast}\coloneqq\frac{\lambda_{\min}}{2\lambda_{\max}^{2}},\qquad\beta_{\ast}\coloneqq\left(1-\frac{\lambda_{\min}^{2}}{4\lambda_{\max}^{2}}\right)^{1/2}.$ (3.14)
This choice of parameters is optimal in the sense that they minimize $\beta$ in (3.13).
Furthermore, by the definition of the weak solution to the Dirichlet problem in (2.1), the $H^{1}(\Omega)$-norm of $u_{D}^{*}$ can be bounded in terms of $\|f\|_{H^{-1}(\Omega)}$ as follows:
$\begin{split}\lambda_{\min}\|u_{D}^{*}\|_{H^{1}(\Omega)}^{2}&\leq\int_{\Omega}\nabla u_{D}^{*}\cdot A\nabla u_{D}^{*}\,\mathrm{d}x+\int_{\Omega}c|u_{D}^{*}|^{2}\,\mathrm{d}x\\ &=\int_{\Omega}fu_{D}^{*}\,\mathrm{d}x\leq\|f\|_{H^{-1}(\Omega)}\|u_{D}^{*}\|_{H^{1}(\Omega)}.\end{split}$ (3.15)
It follows that
$\|u_{D}^{*}\|_{H^{1}(\Omega)}\leq\frac{\|f\|_{H^{-1}(\Omega)}}{\lambda_{\min}}.$ (3.16)
Combining Theorem 3.2 with the bound in (3.16), we obtain
$\|u_{t}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\frac{\|f\|_{H^{-1}(\Omega)}}{\lambda_{\min}}\beta_{\ast}^{t}\quad\text{for all }t\geq 0.$ (3.17)
An analogous estimate holds for the weak solution $u_{N}^{*}$ to the Neumann problem under the Sobolev gradient flow (3.7):
$\|u_{t}-u_{N}^{*}\|_{H^{1}(\Omega)}\leq\frac{\|f\|_{H^{-1}(\Omega)}}{\lambda_{\min}}\beta_{\ast}^{t}\quad\text{for all }t\geq 0.$ (3.18)
Estimates (3.17) and (3.18) yield the following corollary.
Corollary 3.3.
Assume Assumption 2.1 holds. For any $\varepsilon\in(0,1/\lambda_{\min})$, let $\beta_{\ast}$ be defined as in (3.14). If
$T\geq\left\lceil\frac{\log\|f\|_{H^{-1}(\Omega)}+|\log(\varepsilon\lambda_{\min})|}{|\log\beta_{\ast}|}\right\rceil,$
then the iterates $u_{T}$ produced by both schemes (3.6) and (3.7) with stepsize $\alpha^{*}$ in (3.14) satisfy
$\|u_{T}-u_{D}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon,\quad\|u_{T}-u_{N}^{*}\|_{H^{1}(\Omega)}\leq\varepsilon.$
Remark 3.4.
In the estimate (3.15), we use the $H^{-1}(\Omega)$-norm of $f$ instead of the $L^{2}(\Omega)$-norm, since the $H^{-1}$-norm is smaller. This follows from the fact that the Poincaré constant $C_{P}$ for the unit hypercube $\Omega$ is $1/\pi\sqrt{d}$ (see Appendix A.2). That is,
$\|\phi\|_{L^{2}(\Omega)}\leq\frac{1}{\pi\sqrt{d}}\|\nabla\phi\|_{L^{2}(\Omega)}\quad\text{for all }\phi\in H_{0}^{1}(\Omega).$
Consequently, for every $f\in L^{2}(\Omega)$,
$\|f\|_{H^{-1}(\Omega)}$ $\coloneqq\sup\left\{\int_{\Omega}f\phi\,{\,\mathrm{d}}x:\phi\in H_{0}^{1}(\Omega),\;\|\nabla\phi\|_{L^{2}(\Omega)}\leq 1\right\}$
$\leq\sup\left\{\|f\|_{L^{2}(\Omega)}\|\phi\|_{L^{2}(\Omega)}:\phi\in H_{0}^{1}(\Omega),\;\|\nabla\phi\|_{L^{2}(\Omega)}\leq 1\right\}$
$\leq\sup\left\{\|f\|_{L^{2}(\Omega)}C_{P}\|\nabla\phi\|_{L^{2}(\Omega)}:\phi\in H_{0}^{1}(\Omega),\;\|\nabla\phi\|_{L^{2}(\Omega)}\leq 1\right\}$
$\leq\frac{1}{\pi\sqrt{d}}\|f\|_{L^{2}(\Omega)}.$
3.2 Step 2: Barron Norm Estimates
In this section, we focus on proving the following theorem.
Theorem 3.5.
Let $\alpha_{\ast}$ be defined as in (3.14), and set
$p_{D}(d)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{2+\pi}{2\pi}\alpha_{\ast}\ell_{A,D}d^{2}+\alpha_{\ast}\ell_{c}+1,\quad p_{N}(d)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{2+\pi}{2\pi}\alpha_{\ast}\ell_{A,N}d^{2}+\alpha_{\ast}\ell_{c}+1.$ (3.19)
  1. (1)
    Under Assumptions 2.1 and 2.5, the iterative scheme (3.6) for the Dirichlet problem (1.3) satisfies
    $\|u_{t+1}\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq p_{D}(d)\,\|u_{t}\|_{\mathcal{B}_{s}^{2}(\Omega)}+\frac{1}{2}\alpha_{\ast}\ell_{f,D}.$
  2. (2)
    Under Assumptions 2.1 and 2.6, the iterative scheme (3.7) for the Neumann problem (1.4) satisfies
    $\|u_{t+1}\|_{\mathcal{B}_{c}^{2}(\Omega)}\leq p_{N}(d)\,\|u_{t}\|_{\mathcal{B}_{c}^{2}(\Omega)}+\frac{1}{2}\alpha_{\ast}\ell_{f,N}.$
3.2.1 Sobolev Gradient Flow via the Inverse Elliptic Operator
For later use, we define the inverse elliptic operator $(I-\Delta)^{-1}$ under both Dirichlet and Neumann boundary conditions. Given any $\varphi\in L^{2}(\Omega)$, we define:
  • $(I-\Delta)_{D}^{-1}\varphi\in H_{0}^{1}(\Omega)$ denote the unique weak solution to
    $(I-\Delta)w=\varphi\quad\text{in }\Omega,\quad w=0\quad\text{on }\partial\Omega,$
  • $(I-\Delta)_{N}^{-1}\varphi\in H^{1}(\Omega)$ denote the unique weak solution to
    $(I-\Delta)w=\varphi\quad\text{in }\Omega,\quad\frac{\partial w}{\partial\nu}=0\quad\text{on }\partial\Omega.$
The existence and uniqueness of the weak solution in both cases are guaranteed by the Lax–Milgram theorem.
According to (3.3), we have
$\int_{\Omega}D_{H_{0}^{1}}\mathcal{E}_{D}(u)\,v+\nabla D_{H_{0}^{1}}\mathcal{E}_{D}(u)\cdot\nabla v\,\mathrm{d}x=\int_{\Omega}\nabla u\cdot A\nabla v+cuv-fv\,\mathrm{d}x,\quad\text{for all }v\in H_{0}^{1}(\Omega).$
Thus, $D_{H_{0}^{1}}\mathcal{E}_{D}(u)\in H_{0}^{1}(\Omega)$ is the unique weak solution to the elliptic problem
$(I-\Delta)w=\mathcal{L}u-f\quad\text{in }\Omega,\quad w=0\quad\text{on }\partial\Omega,$
where
$\mathcal{L}u\coloneqq-\nabla\cdot A\nabla u+cu$
is the second-order elliptic operator. This implies that the energy gradient for the Dirichlet problem can be expressed as
$D_{H_{0}^{1}}\mathcal{E}_{D}(u)=(I-\Delta)_{D}^{-1}(\mathcal{L}u-f).$
Hence, the iterative scheme (3.6) for the Dirichlet problem (1.3) becomes
$u_{t+1}=u_{t}-\alpha_{\ast}(I-\Delta)^{-1}_{D}(\mathcal{L}u_{t}-f).$ (3.20)
Similarly, from (3.4), we know that $D_{H^{1}}\mathcal{E}_{N}(u)\in H^{1}(\Omega)$ is the unique weak solution to
$(I-\Delta)w=\mathcal{L}u-f\quad\text{in }\Omega,\quad\frac{\partial w}{\partial\nu}=0\quad\text{on }\partial\Omega.$
Therefore, the energy gradient for the Neumann problem is given by
$D_{H^{1}}\mathcal{E}_{N}(u)=(I-\Delta)_{N}^{-1}(\mathcal{L}u-f).$
Hence, the iterative scheme (3.7) for the Neumann problem (1.4) becomes
$u_{t+1}=u_{t}-\alpha_{\ast}(I-\Delta)_{N}^{-1}(\mathcal{L}u_{t}-f).$ (3.21)
3.2.2 Analysis of the Barron Norm
A key ingredient in our analysis is the spectral property of the inverse elliptic operator. The result stated below will be used repeatedly in this section, and its proof follows from a direct computation.
Proposition 3.6.
For each $k\in\mathbb{N}^{d}$, we have
$(I-\Delta)_{D}^{-1}S_{k}(x)=\frac{1}{1+\pi^{2}\|k\|^{2}}S_{k}(x),\quad(I-\Delta)_{N}^{-1}C_{k}(x)=\frac{1}{1+\pi^{2}\|k\|^{2}}C_{k}(x).$
We next collect several algebraic and functional properties of Barron spaces that will be used in the proof of Theorem 3.5.
Proposition 3.7.
Let $\widetilde{\mathcal{B}}^{n}(\Omega)$ denote any of the spaces $\mathcal{B}_{e}^{n}(\Omega)$, $\mathcal{B}_{s}^{n}(\Omega)$, $\mathcal{B}_{c}^{n}(\Omega)$, or $\mathcal{B}_{i,j}^{n}(\Omega)$. Then:
  1. (1)
    (Linearity) For all $g_{1},g_{2}\in\widetilde{\mathcal{B}}^{n}(\Omega)$, and $\lambda_{1},\lambda_{2}\in\mathbb{R}$,
    $\|\lambda_{1}g_{1}+\lambda_{2}g_{2}\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}\leq|\lambda_{1}|\|g_{1}\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}+|\lambda_{2}|\|g_{2}\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}.$
  2. (2)
    (Monotonicity) Let $n\geq m\geq 0$. If $g\in\widetilde{\mathcal{B}}^{n}(\Omega)$, then
    $\|g\|_{\widetilde{\mathcal{B}}^{m}(\Omega)}\leq\|g\|_{\widetilde{\mathcal{B}}^{n}(\Omega)}\quad\text{and}\quad\widetilde{\mathcal{B}}^{n}(\Omega)\subseteq\widetilde{\mathcal{B}}^{m}(\Omega).$
  3. (3)
    (Product estimates)
[…]
  4. (4)
    (Inverse operator estimates)
    $\left\|(I-\Delta)_{D}^{-1}g\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\frac{1}{2}\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}\quad\text{for }g\in\mathcal{B}_{s}^{0}(\Omega),$
    $\left\|(I-\Delta)_{N}^{-1}g\right\|_{\mathcal{B}_{c}^{2}(\Omega)}$ $=\frac{1}{2}\|g\|_{\mathcal{B}_{c}^{0}(\Omega)}\quad\text{for }g\in\mathcal{B}_{c}^{0}(\Omega).$
Proof.
Statements (1) and (2) follow directly from the definitions.
(3). We prove only the first inequality, as the second one is analogous. By the standard uniform convergence argument in analysis, the interchange of summation when taking products in the following proofs is justified, and the details will be omitted.
Let
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x),\quad h(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}C_{k^{\prime}}(x).$
We compute
$g(x)h(x)$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\prod_{j=1}^{d}\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\cdot\frac{1}{2^{d}}\prod_{j=1}^{d}\left(\sin\left(\pi(k_{j}+k_{j}^{\prime})x_{j}\right)+\sin\left(\pi(k_{j}-k_{j}^{\prime})x_{j}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{1}{2^{d}}a_{k}b_{k^{\prime}}S_{k+\epsilon\circ k^{\prime}}(x).$
We estimate
$1+\pi^{n}\|k+\epsilon\circ k^{\prime}\|^{n}\leq 1+\pi^{n}\left(\|k\|+\|k^{\prime}\|\right)^{n}\leq\left(1+\pi^{n}\|k\|^{n}\right)\left(1+\pi^{n}\|k^{\prime}\|^{n}\right).$
Thus, we obtain
$\|gh\|_{\mathcal{B}_{s}^{n}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{1}{2^{d}}a_{k}b_{k^{\prime}}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{n}(\Omega)}$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{1}{2^{d}}|a_{k}b_{k^{\prime}}|\left(1+\pi^{n}\|k+\epsilon\circ k^{\prime}\|^{n}\right)$
$\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}|a_{k}||b_{k^{\prime}}|\left(1+\pi^{n}\|k\|^{n}\right)\left(1+\pi^{n}\|k^{\prime}\|^{n}\right)$
$=\|g\|_{\mathcal{B}_{s}^{n}(\Omega)}\|h\|_{\mathcal{B}_{c}^{n}(\Omega)}.$
(4). We prove only the first identity, as the second follows analogously. Some standard arguments imply that $(I-\Delta)_{D}^{-1}$ and the summation are interchangeable, i.e.,
$(I-\Delta)_{D}^{-1}g=\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x).$
It now follows from Proposition 3.6 and the definition of the $s$-Barron norm that
$\left\|(I-\Delta)_{D}^{-1}g\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}a_{k}(I-\Delta)_{D}^{-1}S_{k}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}a_{k}\frac{1}{1+\pi^{2}\|k\|^{2}}S_{k}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}=\sum_{k\in\mathbb{N}^{d}}|a_{k}|=\frac{1}{2}\|g\|_{\mathcal{B}_{s}^{0}(\Omega)}.$
Lemma 3.8.
Suppose the condition on $A(x)$ in Assumption 2.5 holds, and let $u\in\mathcal{B}_{s}^{2}(\Omega)$. Then for every $i,j=1,\dots,d$, we have
$\left\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ij}\partial_{j}u)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $\leq\frac{1}{\pi}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)},$ (3.22)
$\left\|(I-\Delta)^{-1}_{D}(A_{ij}\partial_{ij}u)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $\leq\frac{1}{2}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}.$ (3.23)
Under Assumption 2.6 and for $u\in\mathcal{B}_{c}^{2}(\Omega)$, the same estimates hold with $(I-\Delta)^{-1}_{N}$, $\ell_{A,N}$, and $\mathcal{B}_{c}^{2}(\Omega)$.
Proof.
We prove the case under Assumption 2.5; the case under Assumption 2.6 follows by the same argument. By the standard uniform convergence argument in analysis, the interchange of the order of summation and differentiation, as well as the interchange of summation when taking products in the following proofs, is justified. We will omit the details when applying these arguments.
Proof of (3.22) in the case $i=j$. Let
$A_{ii}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}C_{k}(x),\quad u(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}S_{k^{\prime}}(x).$
We compute
$\partial_{i}A_{ii}\partial_{i}u$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\left(\partial_{i}C_{k}\partial_{i}S_{k^{\prime}}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\qquad\cdot\sin\left(\pi k_{i}x_{i}\right)\cos\left(\pi k_{i}^{\prime}x_{i}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\cdot\frac{1}{2^{d}}\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i\end{subarray}}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)-\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$\qquad\cdot\left(\sin\left(\pi(k_{i}+k_{i}^{\prime})x_{i}\right)+\sin\left(\pi(k_{i}-k_{i}^{\prime})x_{i}\right)\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\epsilon_{i}\prod_{m=1}^{d}\epsilon_{m}\cdot S_{k+\epsilon\circ k^{\prime}}(x),$
Thus, we obtain
$\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ii}\partial_{i}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\epsilon_{i}\prod_{m=1}^{d}\epsilon_{m}\cdot(I-\Delta)^{-1}_{D}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{i}^{\prime}\epsilon_{i}\prod_{m=1}^{d}\epsilon_{m}\cdot\frac{1}{1+\pi^{2}\|k+\epsilon\circ k^{\prime}\|^{2}}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}k_{i}^{\prime}$
$\leq$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k\|\|k^{\prime}\|^{2}$
$\leq$ $\frac{1}{\pi}\|A_{ii}\|_{\mathcal{B}_{c}^{1}(\Omega)}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\frac{1}{\pi}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}.$
Proof of (3.22) in the case $i\neq j$. Suppose
$A_{ij}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}M_{k,i,j}(x),\quad u(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}S_{k^{\prime}}(x).$
We compute
$\partial_{i}A_{ij}\partial_{j}u$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\qquad\cdot\cos\left(\pi k_{i}x_{i}\right)\sin\left(\pi k_{i}^{\prime}x_{i}\right)\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\cdot\frac{1}{2^{d}}\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq j\end{subarray}}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)-\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$\qquad\cdot\left(\sin\left(\pi(k_{j}+k_{j}^{\prime})x_{j}\right)+\sin\left(\pi(k_{j}-k_{j}^{\prime})x_{j}\right)\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot S_{k+\epsilon\circ k^{\prime}}(x),$
Thus, we obtain
$\|(I-\Delta)^{-1}_{D}(\partial_{i}A_{ij}\partial_{j}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot(I-\Delta)^{-1}_{D}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}k_{j}^{\prime}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot\frac{1}{1+\pi^{2}\|k+\epsilon\circ k^{\prime}\|^{2}}S_{k+\epsilon\circ k^{\prime}}(x)\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}k_{j}^{\prime}$
$\leq$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k\|\|k^{\prime}\|^{2}$
$\leq$ $\frac{1}{\pi}\|A_{ij}\|_{\mathcal{B}_{i,j}^{1}(\Omega)}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\frac{1}{\pi}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}.$
Proof of (3.23) in the case $i=j$. Let
$A_{ii}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}C_{k}(x),\quad u(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}S_{k^{\prime}}(x).$
We compute
$A_{ii}\partial_{ii}u$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}a_{k}b_{k^{\prime}}\left(C_{k}\partial_{ii}S_{k^{\prime}}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\prod_{m=1}^{d}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}-\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\cdot\frac{1}{2^{d}}\prod_{m=1}^{d}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)-\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\prod_{m=1}^{d}\epsilon_{m}\cdot S_{k+\epsilon\circ k^{\prime}}(x),$
Thus, we obtain
$\|(I-\Delta)^{-1}_{D}(A_{ii}\partial_{ii}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\prod_{m=1}^{d}\epsilon_{m}\cdot(I-\Delta)^{-1}_{D}S_{k+\epsilon\circ k^{\prime}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}-\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime 2}\prod_{m=1}^{d}\epsilon_{m}\cdot\frac{1}{1+\pi^{2}\|k+\epsilon\circ k^{\prime}\|^{2}}S_{k+\epsilon\circ k^{\prime}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}^{\prime 2}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k^{\prime}\|^{2}$
$\leq$ $\frac{1}{2}\|A_{ii}\|_{\mathcal{B}_{c}^{0}(\Omega)}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\frac{1}{2}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}.$
Proof of (3.23) in the case $i\neq j$. Suppose
$A_{ij}(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}M_{k,i,j}(x),\quad u(x)=\sum_{k^{\prime}\in\mathbb{N}^{d}}b_{k^{\prime}}S_{k^{\prime}}(x).$
We compute
$A_{ij}\partial_{ij}u$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\qquad\cdot\sin\left(\pi k_{i}x_{i}\right)\cos\left(\pi k_{i}^{\prime}x_{i}\right)\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\cdot\frac{1}{2^{d}}\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)-\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$\quad\cdot\left(\sin(\pi(k_{i}+k_{i}^{\prime})x_{i})+\sin(\pi(k_{i}-k_{i}^{\prime})x_{i})\right)\left(\sin(\pi(k_{j}+k_{j}^{\prime})x_{j})+\sin(\pi(k_{j}-k_{j}^{\prime})x_{j})\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\epsilon_{i}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot S_{k+\epsilon\circ k^{\prime}}(x),$
Thus, we obtain
$\|(I-\Delta)^{-1}_{D}(A_{ij}\partial_{ij}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\epsilon_{i}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot(I-\Delta)^{-1}_{D}S_{k+\epsilon\circ k^{\prime}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\epsilon_{i}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot\frac{1}{1+\pi^{2}\|k+\epsilon\circ k^{\prime}\|^{2}}S_{k+\epsilon\circ k^{\prime}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}b_{k^{\prime}}|k_{i}^{\prime}k_{j}^{\prime}\leq\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}|a_{k}||b_{k^{\prime}}|\|k^{\prime}\|^{2}$
$\leq$ $\frac{1}{2}\|A_{ij}\|_{\mathcal{B}_{i,j}^{0}(\Omega)}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\frac{1}{2}\ell_{A,D}\|u\|_{\mathcal{B}_{s}^{2}(\Omega)}$
Lemma 3.9.
  1. (1)
    Suppose the condition on $c(x)$ in Assumption 2.5 holds, and let $u\in\mathcal{B}_{s}^{2}(\Omega)$. Then we have (I-Δ)^-1_D(cu)_B_s^2(Ω)≤ℓ_cu_B_s^2(Ω).
  2. (2)
    Suppose the condition on $c(x)$ in Assumption 2.6 holds, and let $u\in{\mathcal{B}}_{c}^{2}(\Omega)$. Then we have (I-Δ)^-1_N(cu)_B_c^2(Ω)≤ℓ_cu_B_c^2(Ω).
Proof.
We provide the proof of (1), as the proof of (2) follows by the same argument. By part (3) of Proposition 3.7, we have $cu\in\mathcal{B}_{s}^{2}(\Omega)$. It then follows from part (4) of the same proposition that
$\|(I-\Delta)^{-1}_{D}(cu)\|_{\mathcal{B}_{s}^{2}(\Omega)}\leq\|cu\|_{\mathcal{B}_{s}^{2}(\Omega)}.$
Applying part (3) again yields the desired result. ∎
Proof of Theorem 3.5.
Note that
$\mathcal{L}u-f=-\nabla\cdot A\nabla u+cu-f=-\sum_{i,j=1}^{d}\partial_{i}A_{ij}\partial_{j}u-\sum_{i,j=1}^{d}A_{ij}\partial_{ij}u+cu-f.$
The result follows from the linearity of $(I-\Delta)^{-1}_{D}$ and $(I-\Delta)^{-1}_{N}$, together with (3.20), (3.21), Proposition 3.7, and Lemmas 3.8 and 3.9. ∎
3.3 Step 3: Neural Network Approximation
The following approximation theorem for Barron functions by neural networks is analogous to [EMW22, Theorem 4], and our proof strategy is a refinement of their argument. To the best of our knowledge, this remains the standard approach in the literature; similar strategies are employed in the approximation results for Barron functions in [CLL21, CLLZ23, FL25].
Theorem 3.10.
Let $g\in\mathcal{B}_{e}^{2}(\Omega)$ be an $e$-Barron function. Suppose that the activation function $\sigma$ in the two-layer neural network of the form (1.2) is the cosine function. Then, for any fixed positive integer $k$, there exists a set of parameters
$\left\{(a_{i},w_{i},b_{i})\in\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right\}_{1\leq i\leq k}$
such that
$\left\|\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma\left(w_{i}^{\top}x+b_{i}\right)-g(x)\right\|_{H^{1}(\Omega)}\leq\frac{\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}}{\sqrt{k}}.$ (3.24)
Proof.
Let $\varepsilon>0$ be fixed. Suppose
$g(x)=\sum_{\omega\in\mathbb{Z}^{d}}c_{\omega}e^{i\pi\omega^{\top}x}=\sum_{\omega\in\mathbb{Z}^{d}}c_{\omega}\cos\left(\pi\omega^{\top}x\right)+ic_{\omega}\sin\left(\pi\omega^{\top}x\right)$
satisfies
$\sum_{\omega\in\mathbb{Z}^{d}}|c_{\omega}|\left(1+\pi^{2}\|\omega\|^{2}\right)\leq\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}+\varepsilon.$ (3.25)
We write $c_{\omega}=c_{\omega,R}+ic_{\omega,I}$ with $c_{\omega,R},c_{\omega,I}\in\mathbb{R}$. Since $g(x)$ is real-valued, we have
$g(x)=\operatorname{Re}(g(x))=\sum_{\omega\in\mathbb{Z}^{d}}-c_{\omega,I}\sin\left(\pi\omega^{\top}x\right)+c_{\omega,R}\cos\left(\pi\omega^{\top}x\right).$ (3.26)
Let $\operatorname{sgn}(x)$ denote the sign function, which equals $1$ if $x>0$, 0 if $x=0$, and 1-1 if $x<0$. Define
$\theta_{\omega}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\begin{cases}\arctan\left(c_{\omega,I}/c_{\omega,R}\right),&\text{if }c_{\omega,R}\neq 0,\\ \operatorname{sgn}(c_{\omega,I}){\pi}/2,&\text{if }c_{\omega,R}=0.\end{cases}$
Then, using the standard cosine addition identity, we obtain
$c_{\omega,I}\sin\left(\pi\omega^{\top}x\right)+c_{\omega,R}\cos\left(\pi\omega^{\top}x\right)=|c_{\omega}|\cos\left(\pi\omega^{\top}x+\theta_{\omega}\right),$ (3.27)
where $|c_{\omega}|=\sqrt{c_{\omega,R}^{2}+c_{\omega,I}^{2}}$. Combining (3.27) with (3.26), we obtain
$g(x)=\sum_{\omega\in\mathbb{Z}^{d}}|c_{\omega}|\cos\left(\pi\omega^{\top}x+\theta_{\omega}\right).$ (3.28)
By (3.25), we know that
$Z\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\sum_{\omega\in\mathbb{Z}^{d}}|c_{\omega}|\leq\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}+\varepsilon<\infty.$
Thus, the measure
$\mu(a,w,b)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\sum_{\omega\in\mathbb{Z}^{d}}\frac{|c_{\omega}|}{Z}\delta_{(Z,\pi\omega,\theta_{\omega})}$
defines a probability measure on $\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}$, where $\delta_{(Z,\pi\omega,\theta_{\omega})}$ denotes the Dirac measure on $\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}$. We equip the space $(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}$, endowed with its Borel $\sigma$-algebra, with the product probability measure ${\mu(a,w,b)}^{\otimes k}$. Let
$\Theta\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\left\{(a_{i},w_{i},b_{i})\right\}_{1\leq i\leq k}\in(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k},$
and define the pointwise approximation error between the neural network and the target function $g$ by
$R(\Theta;x)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma\left(w_{i}^{\top}x+b_{i}\right)-g(x),$
which is a random variable on the probability space $\left((\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k},{\mu(a,w,b)}^{\otimes k}\right)$. Then to prove the theorem, it suffices to establish that
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{H^{1}(\Omega)}\leq\frac{\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}^{2}}{k}.$ (3.29)
Observe that the expected $H^{1}$-error decomposes as
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{H^{1}(\Omega)}=\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}+\mathbb{E}_{\mu^{\otimes k}}\|\nabla R(\Theta;x)\|_{L^{2}(\Omega)}^{2},$ (3.30)
We proceed to estimate the two terms on the right-hand side separately.
Expectation of $\|R(\Theta;x)\|_{L^{2}(\Omega)}^{2}$. Since we assume that $\sigma$ is the cosine function, it follows from (3.28) that
$\mathbb{E}_{\mu(a,w,b)}\left[a\sigma(w^{\top}x+b)\right]$ $=\int_{\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}}a\cos\left(w^{\top}x+b\right){\,\mathrm{d}}\mu(a,w,b)$
$=\sum_{\omega\in\mathbb{Z}^{d}}\frac{|c_{\omega}|}{Z}Z\cos\left(\pi\omega^{\top}x+\theta_{\omega}\right)=g(x).$ (3.31)
It follows from (3.31) that for every $1\leq i\neq j\leq k$, we have
$\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)\left(a_{j}\sigma(w_{j}^{\top}x+b_{j})-g(x)\right){\,\mathrm{d}}{\mu(a,w,b)}^{\otimes k}=0.$ (3.32)
Thus, combining (3.32) and (3.31), we compute that
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$
$=$ $\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\int_{\Omega}\left(\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)^{2}{\,\mathrm{d}}x\mathrm{d}{\mu(a,w,b)}^{\otimes k}$
$=$ $\frac{1}{k^{2}}\int_{\Omega}\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(\sum_{i=1}^{k}\left(a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)\right)^{2}{\,\mathrm{d}}{\mu(a,w,b)}^{\otimes k}\mathrm{d}x$
$=$ $\frac{1}{k^{2}}\int_{\Omega}\sum_{i=1}^{k}\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)^{2}{\,\mathrm{d}}{\mu(a,w,b)}^{\otimes k}\mathrm{d}x$
$=$ $\frac{1}{k}\int_{\Omega}\int_{\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}}\left(a\sigma(w^{\top}x+b)-g(x)\right)^{2}{\,\mathrm{d}}\mu(a,w,b)\mathrm{d}x$
$=$ $\frac{1}{k}\int_{\Omega}\int_{\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}}\left(a\sigma(w^{\top}x+b)-\mathbb{E}_{\mu(a,w,b)}\left[a\sigma(w^{\top}x+b)\right]\right)^{2}{\,\mathrm{d}}\mu(a,w,b)\mathrm{d}x$
$=$ $\frac{1}{k}\int_{\Omega}\operatorname{Var}_{\mu(a,w,b)}\left(a\sigma(w^{\top}x+b)\right){\,\mathrm{d}}x.$
Since $\sigma$ is the cosine function, we have
$\operatorname{Var}_{\mu(a,w,b)}\left(a\sigma(w^{\top}x+b)\right)\leq\mathbb{E}_{\mu(a,w,b)}\left[\left(a\sigma(w^{\top}x+b)\right)^{2}\right]\leq\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\right].$
Thus,
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}\leq\frac{m(\Omega)}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\right]=\frac{1}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\right].$ (3.33)
where $m(\Omega)$ denotes the Lebesgue measure of $\Omega$, which equals $1$ since $\Omega=(0,1)^{d}$.
Expectation of $\|\nabla R(\Theta;x)\|_{L^{2}(\Omega)}^{2}$. Let $\langle w_{i},e_{j}\rangle$ denote the standard Euclidean inner product of $w_{i}$ and $e_{j}$ in $\mathbb{R}^{d}$; that is, $\langle w_{i},e_{j}\rangle$ is the $j$-th component of $w_{i}\in\mathbb{R}^{d}$. Differentiating both sides of (3.31) with respect to $x_{j}$, it is straightforward to verify that for every $1\leq i\neq\ell\leq k$ and for each $1\leq j\leq d$, we have
$\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(a_{i}\langle w_{i},e_{j}\rangle\sigma^{\prime}(w_{i}^{\top}x+b_{i})-\partial_{j}g(x)\right)$
$\qquad\qquad\qquad\qquad\cdot\left(a_{\ell}\langle w_{\ell},e_{j}\rangle\sigma^{\prime}(w_{\ell}^{\top}x+b_{\ell})-\partial_{j}g(x)\right){\,\mathrm{d}}\mu^{\otimes k}=0,$
and
$\mathbb{E}_{(a,w,b)}\left[a\left\langle w,e_{j}\right\rangle\sigma^{\prime}(w^{\top}x+b)\right]=\partial_{j}g(x).$
Thus, for every $1\leq j\leq d$, we have
$\mathbb{E}_{\mu^{\otimes k}}\left\|\partial_{j}R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$
$=$ $\int_{\left(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right)^{k}}\int_{\Omega}\left(\partial_{j}\left(\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma\left(w_{i}^{\top}x+b_{i}\right)\right)-\partial_{j}g(x)\right)^{2}{\,\mathrm{d}}x\mathrm{d}\mu^{\otimes k}$
$=$ $\frac{1}{k^{2}}\int_{\Omega}\int_{\left(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right)^{k}}\left(\sum_{i=1}^{k}\left(a_{i}\left\langle w_{i},e_{j}\right\rangle\sigma^{\prime}\left(w_{i}^{\top}x+b_{i}\right)-\partial_{j}g(x)\right)\right)^{2}{\,\mathrm{d}}\mu^{\otimes k}\mathrm{d}x$
$=$ $\frac{1}{k^{2}}\int_{\Omega}\sum_{i=1}^{k}\int_{\left(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right)^{k}}\left(a_{i}\left\langle w_{i},e_{j}\right\rangle\sigma^{\prime}\left(w_{i}^{\top}x+b_{i}\right)-\partial_{j}g(x)\right)^{2}{\,\mathrm{d}}\mu^{\otimes k}\mathrm{d}x$
$=$ $\frac{1}{k}\int_{\Omega}\int_{\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}}\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)-\partial_{j}g(x)\right)^{2}{\,\mathrm{d}}\mu(a,w,b)\mathrm{d}x$
$=$ $\frac{1}{k}\int_{\Omega}\operatorname{Var}_{\mu(a,w,b)}\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)\right){\,\mathrm{d}}x$
$\leq$ $\frac{1}{k}\int_{\Omega}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)\right)^{2}\right]{\,\mathrm{d}}x.$
Since $\sigma$ is the cosine function, we have
$\begin{split}&\mathbb{E}_{\mu^{\otimes k}}\|\nabla R(\Theta;x)\|_{L^{2}(\Omega)}^{2}=\mathbb{E}_{\mu^{\otimes k}}\sum_{j=1}^{d}\left\|\partial_{j}R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}\\ \leq&\frac{1}{k}\int_{\Omega}\sum_{j=1}^{d}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)\right)^{2}\right]{\,\mathrm{d}}x\\ \leq&\frac{1}{k}\int_{\Omega}\sum_{j=1}^{d}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\right)^{2}\right]{\,\mathrm{d}}x=\frac{1}{k}\int_{\Omega}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right]{\,\mathrm{d}}x\\ \leq&\frac{m(\Omega)}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right]=\frac{1}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right].\end{split}$
Combining (3.30), (3.33), and (3.3), we obtain
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{H^{1}(\Omega)}$ $\leq\frac{1}{k}\left(\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\right]+\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\|w\|^{2}\right]\right)$
$=\frac{1}{k}\sum_{\omega\in\mathbb{Z}^{d}}\frac{|c_{\omega}|}{Z}Z^{2}\left(1+\pi^{2}\|\omega\|^{2}\right)=\frac{Z}{k}\left(\sum_{\omega\in\mathbb{Z}^{d}}|c_{\omega}|\left(1+\pi^{2}\|\omega\|^{2}\right)\right)$
$\leq\frac{\left(\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}+\varepsilon\right)^{2}}{k}.$
Letting $\varepsilon\to 0$ yields the desired estimate (3.29), which completes the proof of the theorem. ∎
Proof of Theorem 2.7.
The conclusion follows by combining Proposition 2.4, Corollary 3.3, Theorem 3.5, Proposition 3.7, and Theorem 3.10. ∎
In order to prove the approximation result using the $\operatorname{ReLU}$ activation, we state Theorem 3.11, which is a revised version of in [LLW21, Theorem 17] adapted to the setting of our Barron norm definition. The Barron norm in that work is defined in a way similar to our $c$-Barron norm, but it is based on the $\ell^{1}$-norm $\|k\|_{1}$ rather than the Euclidean norm $\|k\|$ used here. Consequently, their Barron space is substantially smaller than ours. Nevertheless, Lemma 18 and Proposition 19 in [LLW21]—which are key ingredients in the proof of Theorem 17 there—can be extended to our setting by adapting their arguments with suitable modifications. We provide the complete proof in Appendix A.3 for completeness.
Theorem 3.11.
Suppose $g\in{\mathcal{B}}_{s}^{2}(\Omega)$ is an $s$-Barron function of weight 2. Then, for any fixed positive integer $k$, there exist a constant $c\in\mathbb{R}$ with $|c|\leq 2\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}$, and a collection of parameters
$\left\{(a_{i},w_{i},b_{i})\in\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right\}_{1\leq i\leq k}$
satisfying
[…]
such that
$\left\|c+\sum_{i=1}^{k}a_{i}\operatorname{ReLU}(w_{i}^{\top}x+b_{i})-g(x)\right\|_{H^{1}(\Omega)}\leq\frac{\sqrt{q(d)}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}}{\sqrt{k}},$
where $q(d)=256d^{2}+128d+4$. An analogous result also holds if ${\mathcal{B}}_{s}^{2}(\Omega)$ is replaced by ${\mathcal{B}}_{c}^{2}(\Omega)$.
Proof of Theorem 2.9.
The result follows by combining Corollary 3.3, Theorem 3.5, Proposition 3.7, and Theorem 3.11. ∎
Appendix A Omitted Proofs in Sections 2 and 3
A.1 Proof of Theorem 2.2
We denote the domain $(-1,1)^{d}\subseteq\mathbb{R}^{d}$ by $\widetilde{\Omega}$.
[…]
$\left\{\prod_{i=1}^{d}\operatorname{sc}(\pi k_{i}x_{i}):k=(k_{1},\ldots,k_{d})\in\mathbb{N}^{d},\ x=(x_{1},\ldots,x_{d})\in\widetilde{\Omega}\right\}$, (A.1)
[…]
Given a function $g\in L^{2}(\Omega)$, we define its coordinatewise odd extension by
$\widetilde{g}(x)\coloneqq\begin{cases}\operatorname{sgn}(x_{1})\cdots\operatorname{sgn}(x_{d})g(|x_{1}|,\dots,|x_{d}|),&\text{if }x_{m}\neq 0\text{ for all }m=1,\dots,d,\\ 0,&\text{otherwise},\end{cases}$
where $\operatorname{sgn}(t)$ denotes the sign function, defined by
$\operatorname{sgn}(t)=\begin{cases}1,&\text{if }t>0,\\ 0,&\text{if }t=0,\\ -1,&\text{if }t<0.\end{cases}$
Since $\widetilde{g}$ is odd in each coordinate, its expansion in (A.1) contains only sine terms:
[…]
Restricting this expansion to $\Omega$ yields a representation of $g$ in terms of the collection in (2). The uniqueness of the expansion is ensured by the orthogonality of distinct functions in this system, and the coefficients are real since
$a_{k}=2^{d}\int_{\Omega}g(x)S_{k}(x)\,\mathrm{d}x\in\mathbb{R}.$
(3). The proof follows the same argument as in (2) by considering the coordinatewise even extension of $g\in L^{2}(\Omega)$, say
$\widetilde{g}(x)\coloneqq\begin{cases}g(|x_{1}|,\dots,|x_{d}|),&\text{if }x_{m}\neq 0\text{ for all }m=1,\dots,d,\\ 0,&\text{otherwise}.\end{cases}$
(4). The proof follows the same argument as in (2) by considering an extension of $g\in L^{2}(\Omega)$ that is odd in the $i$-th and $j$-th coordinates and even in the remaining ones, say
$\widetilde{g}(x)\coloneqq\begin{cases}\operatorname{sgn}(x_{i})\operatorname{sgn}(x_{j})g(|x_{1}|,\dots,|x_{d}|),&\text{if }x_{m}\neq 0\text{ for all }m=1,\dots,d,\\ 0,&\text{otherwise}.\end{cases}$
A.2 The Poincaré Constant for the Unit Hypercube
Theorem A.1.
The Poincaré constant $C_{P}$ for the unit hypercube $\Omega$ is $1/\pi\sqrt{d}$; that is,
$C_{P}=\frac{1}{\pi\sqrt{d}}$
is the optimal constant such that
$\|\phi\|_{L^{2}(\Omega)}\leq C_{P}\|\nabla\phi\|_{L^{2}(\Omega)}\quad\text{for all }\phi\in H_{0}^{1}(\Omega).$
Proof.
It is easy to see that
$\frac{1}{C_{P}}=\inf_{\begin{subarray}{c}\phi\in H_{0}^{1}(\Omega)\\ \phi\neq 0\end{subarray}}\frac{\|\nabla\phi\|_{L^{2}(\Omega)}}{\|\phi\|_{L^{2}(\Omega)}}.$
This quantity is the square root of the smallest eigenvalue of the Dirichlet Laplacian (cf. [Eva10, Theorem 6.5.2]):
$\Delta u=\lambda u\quad\text{in }\Omega,\quad u=0\quad\text{on }\partial\Omega.$ (A.2)
It is well known that, for $\Omega=(0,1)^{d}$, the eigenfunctions of the Dirichlet Laplacian are
$\left\{S_{k}(x)=\prod_{i=1}^{d}\sin(\pi k_{i}x_{i}):k=(k_{1},\ldots,k_{d})\in\mathbb{N}_{+}^{d}\right\},$
and the corresponding eigenvalue for $S_{k}(x)$ is $\pi^{2}\|k\|^{2}$ (cf. [CH89, Chapter 6.4.1]). The smallest eigenvalue of (A.2) is therefore $\pi^{2}d$. Hence,
$C_{P}=\frac{1}{\sqrt{\pi^{2}d}}=\frac{1}{\pi\sqrt{d}}.$
A.3 Proof of Theorem 3.11
As mentioned in Section 3.3, our strategy is a revision and generalization of the methods in [LLW21]. In particular, Lemma A.2 generalizes their Lemma 18, Lemma A.3 generalizes their Proposition 19, and Lemma A.4 is simply a restatement of their Lemma 16.
We let $I_{d}\coloneqq[-\sqrt{d},\sqrt{d}]$ throughout the arguments.
Lemma A.2.
Suppose $g\in C^{2}(I_{d})$ and there exists $B>0$ such that
$\|g^{(s)}\|_{L^{\infty}(I_{d})}\leq B\quad\text{for all }s=0,1,2,$
and there exists $\alpha\in(-\sqrt{d},\sqrt{d})$ such that $g^{\prime}(\alpha)=0$. Then, for any positive integer $m$, there exists a function $g_{m}(z)$ defined on $I_{d}$ of the form
$g_{m}(z)=c+\sum_{i=1}^{2m}a_{i}\operatorname{ReLU}(\varepsilon_{i}z+b_{i}),$ (A.3)
where
$|c|\leq B,\quad|a_{i}|\leq\frac{4\sqrt{d}B}{m},\quad\varepsilon_{i}\in\{\pm 1\},\quad|b_{i}|\leq\sqrt{d},\quad\text{for all }1\leq i\leq k,$ (A.4)
such that
$\|g-g_{m}\|_{H^{1}(I_{d})}\leq\frac{\sqrt{10}d^{5/4}B}{m}.$ (A.5)
Proof.
Let $\{z_{j}\}_{0\leq j\leq 2m}$ be a partition of $I_{d}$ with
$z_{0}=-\sqrt{d},\quad z_{m}=\alpha,\quad z_{2m}=\sqrt{d},$
$z_{j+1}-z_{j}=h_{1}\coloneqq\frac{\alpha+\sqrt{d}}{m},\quad j=0,\ldots,m-1,$
$z_{j+1}-z_{j}=h_{2}\coloneqq\frac{\sqrt{d}-\alpha}{m},\quad j=m,\ldots,2m-1.$
Let $g_{m}(z)$ defined on $I_{d}$ be the piecewise linear interpolation of $g(z)$ with respect to $\{z_{j}\}_{0\leq j\leq 2m}$, i.e.,
$g_{m}(z)\coloneqq\begin{cases}g(z_{j+1})\dfrac{z-z_{j}}{h_{1}}+g(z_{j})\dfrac{z_{j+1}-z}{h_{1}},&z\in[z_{j},z_{j+1}],\ j=0,\ldots,m-1,\\[5.16663pt] g(z_{j+1})\dfrac{z-z_{j}}{h_{2}}+g(z_{j})\dfrac{z_{j+1}-z}{h_{2}},&z\in[z_{j},z_{j+1}],\ j=m,\ldots,2m-1.\end{cases}$
Then, according to [AG11, Chapter 11], and using $h_{1},h_{2}\leq 2\sqrt{d}/m$, we obtain
$\|g-g_{m}\|_{L^{\infty}(I_{d})}\leq\frac{\max\{h_{1},h_{2}\}^{2}}{8}\,\|g^{\prime\prime}\|_{L^{\infty}(I_{d})}\leq\frac{dB}{2m^{2}}\leq\frac{dB}{m}.$ (A.6)
We also claim that
$\|g^{\prime}-g_{m}^{\prime}\|_{L^{\infty}(I_{d})}\leq\frac{2\sqrt{d}B}{m}\leq\frac{2dB}{m}.$ (A.7)
Indeed, if $z\in[z_{j},z_{j+1}]$ for some $0\leq j\leq m-1$, then by the mean value theorem there exist $\xi,\eta\in(z_{j},z_{j+1})$ such that
$|g^{\prime}(z)-g_{m}^{\prime}(z)|$ $=\left|g^{\prime}(z)-\frac{g(z_{j+1})-g(z_{j})}{h_{1}}\right|=|g^{\prime}(z)-g^{\prime}(\xi)|$
$=|g^{\prime\prime}(\eta)||z-\xi|\leq|g^{\prime\prime}(\eta)|h_{1}\leq\frac{2\sqrt{d}B}{m}.$
The same bound follows by the same argument when $z\in[z_{j},z_{j+1}]$ for some $m\leq j\leq 2m-1$. This proves (A.7). The estimate (A.5) then follows by combining (A.6), (A.7), and
$\|g-g_{m}\|_{H^{1}(I_{d})}^{2}$ $=\|g-g_{m}\|_{L^{2}(I_{d})}^{2}+\|g^{\prime}-g_{m}^{\prime}\|_{L^{2}(I_{d})}^{2}$
$\leq 2\sqrt{d}\|g-g_{m}\|_{L^{\infty}(I_{d})}^{2}+2\sqrt{d}\|g^{\prime}-g_{m}^{\prime}\|_{L^{\infty}(I_{d})}^{2}.$
Now we show that $g_{m}(z)$ can be written in the ReLU form (A.3) and satisfies the coefficient bound in (A.4). Indeed, it is straightforward to verify that
$g_{m}(z)=g(z_{m})+\sum_{i=1}^{m}a_{i}\operatorname{ReLU}(-z+z_{i})+\sum_{i=m+1}^{2m}a_{i}\operatorname{ReLU}(z-z_{i-1}),\quad z\in I_{d},$
where
$a_{i}=\begin{cases}\dfrac{g(z_{i-1})-2g(z_{i})+g(z_{i+1})}{h_{1}},&i=1,\ldots,m-1,\\[4.30554pt] \dfrac{g(z_{m-1})-g(z_{m})}{h_{1}},&i=m,\\[4.30554pt] \dfrac{g(z_{m+1})-g(z_{m})}{h_{2}},&i=m+1,\\[4.30554pt] \dfrac{g(z_{i-2})-2g(z_{i-1})+g(z_{i})}{h_{2}},&i=m+2,\ldots,2m.\end{cases}$
It remains to verify that $|a_{i}|$ satisfies the upper bound in (A.4).
Case $i=1,\ldots,m-1$. By the Taylor expansion, there exist $\eta_{1}\in(z_{i-1},z_{i})$ and $\eta_{2}\in(z_{i},z_{i+1})$ such that
$g(z_{i-1})$ $=g(z_{i})-g^{\prime}(z_{i})h_{1}+g^{\prime\prime}(\eta_{1})h_{1}^{2},$
$g(z_{i+1})$ $=g(z_{i})+g^{\prime}(z_{i})h_{1}+g^{\prime\prime}(\eta_{2})h_{1}^{2}.$
Hence,
$|a_{i}|=\left|\frac{g(z_{i-1})-2g(z_{i})+g(z_{i+1})}{h_{1}}\right|=|g^{\prime\prime}(\eta_{1})h_{1}+g^{\prime\prime}(\eta_{2})h_{1}|\leq 2Bh_{1}\leq\frac{4\sqrt{d}B}{m}.$
Case $i=m$. By the construction of the partition, $g^{\prime}(z_{m})=g^{\prime}(\alpha)=0$. Hence, by the mean value theorem there exist $\xi,\eta\in(z_{m-1},z_{m})$ such that
$|a_{m}|=\left|\frac{g(z_{m-1})-g(z_{m})}{h_{1}}\right|=|g^{\prime}(\xi)|=|g^{\prime}(\xi)-g^{\prime}(z_{m})|=|g^{\prime\prime}(\eta)|\,h_{1}\leq\frac{2\sqrt{d}B}{m}.$
Case $i=m+1$. By a similar argument to that in the case $i=m$, we obtain the same bound.
Case $i=m+2,\ldots,2m$. By a similar argument to that in the case $1\leq i\leq m-1$, we obtain the same bound. ∎
Lemma A.3.
For $B>0$, define the following families of functions on $\Omega=(0,1)^{d}\subseteq{\mathbb{R}}^{d}$ (it is easy to verify that they are all in $H^{1}(\Omega)$):
$\mathscr{F}_{\sin}(B)$ $\coloneqq\left\{\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi k^{\top}x):|A|\leq B,\ k\in\mathbb{N}^{d}\setminus\{0\}\right\},$
$\mathscr{F}_{\cos}(B)$ $\coloneqq\left\{\frac{A}{1+\pi^{2}\|k\|^{2}}\cos(\pi k^{\top}x):|A|\leq B,\ k\in\mathbb{N}^{d}\setminus\{0\}\right\},$
$\mathscr{F}_{\operatorname{ReLU}}(B)$ $\coloneqq\left\{c+A\operatorname{ReLU}(w^{\top}x+b):|c|\leq B,\ |A|\leq 8\sqrt{d}B,\ \|w\|=1,\ |b|\leq\sqrt{d}\right\}.$
Then the $H^{1}(\Omega)$-closures of the convex hulls of $\mathscr{F}_{\sin}(B)$ and $\mathscr{F}_{\cos}(B)$ are both contained in the $H^{1}(\Omega)$-closure of the convex hull of $\mathscr{F}_{\operatorname{ReLU}}(B)$.
Proof.
We only prove the sine function case, as the cosine case follows similarly. It suffices to show that
$\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi k^{\top}x)\in\text{the $H^{1}(\Omega)$-closure of }\operatorname{conv}\left(\mathscr{F}_{\operatorname{ReLU}}(B)\right).$
By Lemma A.2, the function $\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi\|k\|z)$ defined on $I_{d}$ can be $H^{1}(I_{d})$-approximated by a linear combination of a constant and terms of the form $\operatorname{ReLU}(\varepsilon z+b)$, with the sum of the absolute values of the coefficients of $\operatorname{ReLU}(\varepsilon z+b)$ bounded by $8\sqrt{d}B$. Hence, $\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi\|k\|z)$ is contained in the $H^{1}(I_{d})$-closure of
$\operatorname{conv}\left\{c+A\operatorname{ReLU}(\varepsilon z+b):|c|\leq B,\ |A|\leq 8\sqrt{d}B,\ \varepsilon\in\{\pm 1\},\ |b|\leq\sqrt{d}\right\}.$
Since $\|(k/\|k\|)^{\top}x\|\leq\|x\|\leq\sqrt{d}$, it follows that
$\frac{A}{1+\pi^{2}\|k\|^{2}}\sin(\pi k^{\top}x)=\frac{A}{1+\pi^{2}\|k\|^{2}}\sin\!\left(\pi\|k\|\,(k/\|k\|)^{\top}x\right)$
also lies in the $H^{1}(\Omega)$-closure of
$\operatorname{conv}\left\{c+A\operatorname{ReLU}\left(\varepsilon(k/\|k\|)^{\top}x+b\right):|c|\leq B,\ |A|\leq 8\sqrt{d}B,\ \varepsilon\in\{\pm 1\},\ |b|\leq\sqrt{d}\right\}.$ (A.8)
The result then follows from the fact that (A.8) is contained in $\operatorname{conv}\left(\mathscr{F}_{\operatorname{ReLU}}(B)\right)$. ∎
Lemma A.4 ([Pis81, Bar93]).
Let $\mathscr{G}$ be a subset of a Hilbert space such that the norm of every element in $\mathscr{G}$ is bounded by $B_{\mathscr{G}}>0$. Suppose $u$ belongs to the closure of the convex hull of $\mathscr{G}$. Then, for every positive integer $k$, there exist $\{g_{i}\}_{i=1}^{k}\subseteq\mathscr{G}$ and $\{c_{i}\}_{i=1}^{k}\subseteq[0,1]$ with $\sum_{i=1}^{k}c_{i}=1$ such that
$\left\|u-\sum_{i=1}^{k}c_{i}g_{i}\right\|\leq\frac{B_{\mathscr{G}}}{\sqrt{k}}.$
Proof of Theorem 3.11.
We only prove the case for ${\mathcal{B}}_{s}^{2}(\Omega)$, as the case for ${\mathcal{B}}_{c}^{2}(\Omega)$ follows similarly.
Although the definition of an $s$-Barron function involves the basis $\{S_{k}(x)\}$ with $k=(k_{1},\ldots,k_{d})\in\mathbb{N}^{d}$ such that all $k_{i}\neq 0$, for the purpose of unifying the proof with the ${\mathcal{B}}_{c}^{2}(\Omega)$ case, we set $a_{k}=0$ for all $k\in\mathbb{N}^{d}$ having at least one component $k_{i}=0$ in the expansion
$g(x)=\sum_{k\in\mathbb{N}^{d}}a_{k}S_{k}(x).$
With this convention, we can rewrite
$g(x)-a_{0}=\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}a_{k}S_{k}(x)=\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}a_{k}\cdot\frac{1}{2^{d}}\sum_{\widetilde{k}\in\{(\pm k_{1},\ldots,\pm k_{d})\}}\varepsilon\operatorname{sc}(\pi\widetilde{k}^{\top}x)$
$=$ $\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}\sum_{\widetilde{k}\in\{(\pm k_{1},\ldots,\pm k_{d})\}}\frac{\varepsilon a_{k}\,(1+\pi^{2}\|\widetilde{k}\|^{2})}{2^{d}A_{g}}\cdot\frac{A_{g}}{1+\pi^{2}\|\widetilde{k}\|^{2}}\,\operatorname{sc}(\pi\widetilde{k}^{\top}x).$
Here $\varepsilon\in\{\pm 1\}$, $\operatorname{sc}$ denotes $\cos$ if $d$ is even and $\sin$ if $d$ is odd, and
$A_{g}\coloneqq\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}|a_{k}|\,\left(1+\pi^{2}\|k\|^{2}\right)\leq\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}.$
(In the ${\mathcal{B}}_{c}^{2}(\Omega)$ case, only the cosine term appears.) Under this normalization, we have
$\sum_{k\in\mathbb{N}^{d}\setminus\{0\}}\sum_{\widetilde{k}\in\{(\pm k_{1},\ldots,\pm k_{d})\}}\left|\frac{\varepsilon a_{k}\,(1+\pi^{2}\|\widetilde{k}\|^{2})}{2^{d}A_{g}}\right|=1,$
Next, we observe that
$\|g\|_{H^{1}(\Omega)}^{2}=\sum_{k\in\mathbb{N}^{d}}|\alpha_{k}|(1+\pi^{2}\|k\|^{2})|a_{k}|^{2}\leq\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}^{2}<\infty,$
where $\alpha_{k}$ is a constant depending on $k$ with $|\alpha_{k}|\leq 1$. This shows that $g(x)\in H^{1}(\Omega)$, and hence $g(x)-a_{0}\in H^{1}(\Omega)$ as well.
Combining the above observations, we conclude that $g(x)-a_{0}$ lies in the $H^{1}(\Omega)$-closure of either $\operatorname{conv}(\mathscr{F}_{\cos}(\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}))$ or $\operatorname{conv}(\mathscr{F}_{\sin}(\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}))$. Hence, by Lemma A.3, it is contained in the $H^{1}(\Omega)$-closure of $\operatorname{conv}(\mathscr{F}_{\operatorname{ReLU}}(\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}))$. Moreover, since $|a_{0}|\leq\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}$ (not needed here as $a_{0}=0$, but required for the ${\mathcal{B}}_{c}^{2}(\Omega)$ case; we retain it for consistency of the proofs), it follows that $g(x)$ lies in the $H^{1}(\Omega)$-closure of the convex hull of
$\mathscr{G}\coloneqq\Big\{c+A\operatorname{ReLU}(w^{\top}x+b):$
$\qquad\qquad\qquad\qquad|c|\leq 2\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)},\ |A|\leq 8\sqrt{d}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)},\ \|w\|=1,\ |b|\leq\sqrt{d}\Big\}.$
The result then follows from Lemma A.4 together with the bound for the $H^{1}$-norm of each $h\in\mathscr{G}$:
$\|h\|_{H^{1}(\Omega)}^{2}$ $\leq\left(2\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}+8\sqrt{d}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}(\sqrt{d}+\sqrt{d})\right)^{2}+(8\sqrt{d}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)})^{2}$
$=(256d^{2}+128d+4)\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)}^{2}.$
Acknowledgments
Z. Chen thanks Jianfeng Lu and Yulong Lu for helpful discussions. L. Huang gratefully acknowledges Padmavathi Srinivasan for her continuous encouragement and patient guidance during his first two years of graduate study. He also thanks Wenkui Du, Lev Fedorov, Mengxuan Yang, and Shengxuan Zhou for many valuable discussions. We used ChatGPT for minor language editing; all technical content was written and verified by the authors.